### v1

# Sonja Hohloch  &\&&  Pedro Santos

Abstract.
We adapt and generalize McDuff’s method of probes from toric system to so called semitoric integrable systems and apply it to study the (non)displaceability properties of the fibers of $3$ examples of semitoric integrable systems.
MSC codes. Primary: 53D12, 57R17, 70H06. Secondary: 53D40, 53D20.
1. Introduction
Let $(M,\omega)$ be a closed symplectic manifold. The Arnold conjecture, formulated in the $1960$s, can be stated in many different ways: for instance via the fixed points of a Hamiltonian diffeomorphism $\phi$ or equivalently looking at the intersection points of the graph of $\phi$ and the diagonal in $M\times M$. When one endows $M\times M$ with the symplectic form $\omega\oplus-\omega$ the graph of $\phi$ and the diagonal become Lagrangian submanifolds. Therefore the question about the existence of fixed points is turned into a so called Lagrangian intersection problem. This was the approach taken by Floer [floer1988morse, floer1989witten, floer1988unregularized] towards the end of the 1980s to prove the Arnold conjecture under certain assumptions. Roughly speaking, for a Lagrangian submanifold $L\subset M$ with $\pi_{2}(M,L)=\{0\}$, Floer developed a homology theory generated by the intersection points of $L\cap\phi(L)$ where $\phi$ is a Hamiltonian diffeomorphism.
Moreover, at the beginning of the 1980s, Gromov’s non-squeezing theorem promoted an interest in questions of symplectic rigidity. A Lagrangian intersection problem can be seen as an example of a symplectic rigidity phenomenon, since $\phi(L)\cap L\stackrel{{\scriptstyle?}}{{=}}\emptyset$ can be interpreted as if $L$ can be moved away from itself or not.
More formally, given a closed symplectic manifold $(M,\omega)$ with a subset $A$ we say that $A$ is displaceable if there exists a Hamiltonian diffeomorphism $\phi$ such that $A\cap\phi(A)=\emptyset$. Otherwise it is called nondisplaceable.
Therefore, the interest in (non)displaceability is at least twofold: on the one hand motivated by the computation of Lagrangian Floer homology and on the other hand by symplectic rigidity. Over the years, research on (non)displaceability developed into an area of its own, for instance, just to mention a few, it is essential for Fukaya categories (see Fukaya & Oh & Ohta & Ono [fukayabook1, fukayabook2]), it motivated the development of symplectic quasi-states by Entov & Polterovich [polterovich2006quasi], and the development of explicit methods, see for example McDuff [mcduff2011displacing] and Abreu & Borman & McDuff [abreu2014displacing], that give sufficient conditions for a Lagrangian torus fiber to be displaceable. This last aspect is in fact the starting point for the present paper, as we will explain now.
Briefly summarized, McDuff’s [mcduff2011displacing] explicit method (called the method of probes) makes use of the Delzant polytope (= momentum polytope) of a symplectic toric manifold. In the world of integrable systems, toric systems are the simplest examples and are classified by their corresponding Delzant polytope. The next natural, more general class of systems are so called semitoric systems (see Definition 2.17) which were classified by Pelayo & Vũ Ngǫc [pelayo2009semitoric, pelayo2011constructing] and Palmer & Pelayo & Tang [palmer2019semitoric] in terms of $5$ invariants. One of these invariants, the so called polytope invariant, is the generalization of the Delzant polytope of a toric system. In the present paper we will generalize the method of probes from toric system to semitoric systems and apply it to several examples.
1.1. Our results
Let $(M,\omega,F=(J,H))$ be a semitoric system (see Definition 2.17) on a $4$-dimensional symplectic manifold. Intuitively semitoric systems are more general than toric systems by requiring only for an $S^{1}$-symmetry instead of a $\mathbb{T}^{2}$-symmetry and allowing for focus-focus singularities in addition to the ones appearing in the toric case. As in the toric situation the fibers are connected. A fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of Lagrangian $k$ spheres joint at the poles. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded. Our first result is about the (non)displaceability of focus-focus fibers with more than one focus-focus point:
Theorem 1.1.
A focus-focus fiber containing more than one focus-focus point is nondisplaceable.
This is proven in Section 3.2. It is a consequence of the topological nondisplaceability of Lagrangian spheres. An explicit example of a semitoric system containing focus-focus fibers with $2$ pinches can be found in Section 4.3.
Therefore the question of (non)displaceability of focus-focus fibers is reduced to considering fibers containing a single focus-focus point. In Section 3.3 we give conditions, in terms of the invariants of the semitoric system, to displace a focus-focus fiber.
The so called semitoric polytope invariant of $(M,\omega,F)$ arises from $F(M)$ by introducing vertical cuts, from above or below, to the focus-focus values and then "straightening" it to obtain a convex polytope, as illustrated in Figure 1.1. This yields a representative of the polytope invariant, for the detailed procedure see Section 2.8.
Figure 1.1. Straightening homeomorphism $f_{\vec{\epsilon}}$ applied to $F(M)$ to obtain a representative of the polytope invariant $f_{\vec{\epsilon}}(F(M))$. The focus-focus values of $F(M)$ are $c_{1}$ and $c_{2}$.
McDuff’s method of probes first embeds a disk of large enough radius times a circle in an area preserving way in $(M,\omega)$. In the Delzant polytope, this corresponds to choosing a suitable ray that starts from an edge of the polytope. We apply this idea in a representative of the polytope invariant of a semitoric system to displace semitoric fibers. Compared to the toric situation, we have to deal with additional obstructions coming from the presence of focus-focus values: the cuts introduced in $F(M)$ to obtain a representative of the polytope invariant act as barriers for the ray. Also different from the toric situation, in order to displace a focus-focus fiber we displace an open neighborhood of it, which requires "more space", i.e., it creates the need for more conditions. The idea is to symplectically embed the product of two disks such that one disk has big enough radius to be able to apply McDuff’s ideas. In order to do this procedure one takes into consideration the distance to the boundary of the focus-focus value in a representative of the polytope invariant while paying attention to the cuts introduced by the other focus-focus values, see Figure 1.2.
Figure 1.2. $A(f_{\vec{\epsilon}}(F(M)))$ is the result of applying a suitable integral affine transformation $A$ to the representative of the polytope invariant $f_{\vec{\epsilon}}(F(M))$. Let $0<h<a<\frac{R}{2}$. The square $[0,a]\times[0,a]$ fits inside the rectangle $[0,R]\times[0,a]$. The rectangle is inside $A(f_{\vec{\epsilon}}(F(M)))$ and does not intersect the cut associated with the focus-focus value $c_{2}$. The rectangle $[0,R]\times[0,a]$ represents the symplectic embedding of the product of disks $D^{2}(R)\times D^{2}(a)$. Under these conditions the focus-focus fiber given by $c_{1}$ is displaceable, since we can displace $D^{2}(a)\times D^{2}(a)$ inside $D^{2}(R)\times D^{2}(a)$ due to the fact that $a<\frac{R}{2}$ and we can "torically smoothen" the corner introduced by $c_{1}$.
Theorem 1.2.
In a semitoric system a focus-focus fiber can be displaced as long as there exists a symplectic chart with "enough space", in the sense of Figure 1.2, to accommodate the product of the aforementioned disks.
This is the content of Section 3.1 and Section 3.3.
In Section 4 we apply our results to study the (non)displaceability properties of semitoric fibers in $3$ examples:
First, in Section 4.1 we study the coupled spin oscillator which is a semitoric system $(M,\omega,F)$ on $M=S^{2}\times\mathbb{R}^{2}$, for which the main result is:
Theorem 1.3.
All of the fibers of a coupled spin-oscillator are displaceable.
In Section 4.2 we study the coupled angular momenta and its variations. Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=S^{2}\times S^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{S^{2}}\bigoplus R_{2}\omega_{S^{2}})$ where $\omega_{S^{2}}$ is the standard symplectic form on $S^{2}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, with $R_{1}<R_{2}$. The case $R_{1}>R_{2}$ is called the reverse coupled angular momenta and the case $R_{1}=R_{2}$ is called the Kepler problem. For an appropriate range of the parameter $t$ a focus-focus point always exists.
Proposition 1.4.
Let $R_{1}\neq R_{2}$ and $t$ be in the parameter range such that a focus-focus point exists. Then
  • there only exists one nondisplaceable fiber (for which we give an explicit formula in Section 4.2) in the coupled angular momenta system. In particular it is nondisplaceable by symplectomorphisms.
  • All other fibers are displaceable.
  • In particular, the focus-focus fiber is displaceable.
Note that the symplectic rigidity problem is sensitive to the choice of $R_{1}$ and $R_{2}$. For the Kepler problem concerning the question of (non)displaceability we get a different behavior:
Proposition 1.5.
Let $R_{1}=R_{2}$ and $t$ be in the parameter range such that a focus-focus point exists. There exists a $t_{0}$ such that:
  • If $t<t_{0}$ there exists only a single nondisplaceable fiber. In particular the focus-focus fiber is displaceable.
  • If $t=t_{0}$ the focus-focus fiber is nondisplaceable. In particular all other fibers are displaceable, and the focus-focus fiber is nondisplaceable by symplectomorphisms.
  • If $t>t_{0}$ there exist an infinite amount of nondisplaceable fibers. In particular the focus-focus fiber is nondisplaceable.
In Section 4.3 we start studying the (non)displaceability of the fibers of the toric system induced by the octagon $\Delta$ with vertices
$\{(1,0),(0,1),(2,0),(0,2),(1,3),(2,3),(3,1),(3,2)\}$
as a Delzant polytope. Then we consider the semitoric perturbation of this system, introduced in De Meulenaere & Hohloch [de2021family], and study the (non)displaceability properties of the semitoric fibers of this perturbation. Let $(M,\omega,F_{t})$ be the semitoric perturbation of the toric system. For $t=\frac{1}{2}$ the system has two focus-focus fibers with two focus-focus points and otherwise $4$ focus-focus fibers with a single focus-focus point. In Appendix A we show that a representative of the polytope invariant, for $t\neq\frac{1}{2}$, is given by the octagon $\Delta$.
Proposition 1.6.
Let $(M,\omega,F_{t})$ be the semitoric system induced by the octagon. Then
  • For $t=\frac{1}{2}$
    1. (1)
      The focus-focus fibers and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in any representative of the polytope invariant are nondisplaceable.
    2. (2)
      All other fibers are displaceable.
  • For $t\neq\frac{1}{2}$
    1. (1)
      If the image of the focus-focus values in the representative of the polytope invariant given by the octagon are contained in the region $\{y>2\}\cap\{y<1\}$ then the fibers over the points $\{(1,1),(2,1),(\frac{3}{2},\frac{3}{2}),(1,2),(2,2)\}$ are nondisplaceable. All other fibers are displaceable. In particular the focus-focus fibers are displaceable.
    2. (2)
      If the image of the focus-focus values in the representative of the polytope invariant given by the octagon are contained in the region $\{y\leq 2\}\cap\{y\geq 1\}$ then the focus-focus fibers and the fiber over the point $\{(\frac{3}{2},\frac{3}{2})\}$ are nondisplaceable. All other fibers are displaceable.
If one wants to study (non)displaceability questions in integrable systems the next natural step is to consider so called hypersemitoric systems, or even more general integrable systems. This would mean dealing with fibers containing hyperbolic or degenerate singularities which is beyond the scope of this paper.
1.2. Structure of the paper
  • In Section 2 we give the necessary definitions and conventions and recall essential results from the literature.
  • In Section 3 we generalize McDuff’s method of probes and give enough conditions in order to be able to displace focus-focus fibers in semitoric systems. Furthermore, we prove that a focus-focus fiber containing more than a single focus-focus points is nondisplaceable.
  • In Section 4 we study the (non)displaceability properties of $3$ semitoric systems, namely the coupled spin oscillator, the coupled angular momenta and the semitoric system induced by the octagon.
  • In the Appendix A we compute a representative of the polytope invariant of the semitoric system induced by the octagon.
Acknowledgments
The first author was partially and second author fully supported by the FWO-EoS project Beyond symplectic geometry with UA Antigoon number 45816. Moreover, the first author was also partially supported by the grant "Francqui onderzoekshoogleraar 2023-2026" of the Francqui Foundation with UA Antigoon number 49741.
2. Preliminaries and background
2.1. Notions and conventions from symplectic geometry
A symplectic manifold is a pair $(M,\omega)$ where $M$ is a smooth manifold and $\omega$ is a closed non-degenerate $2$-form. A submanifold $Y$ of $M$ is a Lagrangian submanifold if, at each $p\in Y$, the tangent space $T_{p}Y$ is a Lagrangian subspace of $T_{p}M$, i.e., $\omega_{p}|_{T_{p}Y}\equiv 0$ and $\dim(T_{p}Y)=\frac{1}{2}\dim(T_{p}M)$.
Let $H:M\rightarrow\mathbb{R}$ be a smooth function. Due to the non-degeneracy of $\omega$, there exists a unique vector field $X_{H}$ on $M$ such that $\omega(X_{H},\cdot)=dH$. We call $X_{H}$ the Hamiltonian vector field of $H$, and its flow $\phi^{t}_{H}$ the Hamiltonian flow of $H$. The Poisson bracket of two functions $f,g\in C^{\infty}(M,\mathbb{R})$ induced by $\omega$ is
$\{f,g\}:=\omega(X_{f},X_{g}).$
Now let $(M,\omega)$ be a symplectic manifold. We say that a subset $X\subset M$ cannot be displaced from a subset $Y\subset M$ if
$\phi^{1}_{H}(X)\cap\overline{Y}\neq\emptyset,\quad\forall H\in C^{\infty}_{c}(M,\mathbb{R}).$
where $\overline{Y}$ denotes the closure of $Y$. If $X$ cannot be displaced from itself we call it nondisplaceable and otherwise displaceable.
2.2. Integrable systems
[…]
  • $f_{1},...,f_{n}$ are in involution, i.e., $\{f_{i},f_{j}\}=0$ for all $1\leq i,j\leq n$;
  • $f_{1},...,f_{n}$ are independent, i.e., $X_{f_{1}},...,X_{f_{n}}$ are linearly independent almost everywhere on $M$.
We call the map $F:=(f_{1},...,f_{n}):M\rightarrow\mathbb{R}^{n}$ the momentum map of the integrable system.
Two integrable systems $(M,\omega,F)$ and $(M^{\prime},\omega^{\prime},F^{\prime})$ are isomorphic if there exists a diffeomorphism $\psi:M\rightarrow M^{\prime}$ such that $\psi^{*}\omega^{\prime}=\omega$ and a diffeomorphism $g:F(M)\rightarrow F^{\prime}(M^{\prime})$ such that $F^{\prime}\circ f=g\circ F$.
Henceforth we assume $M$ to be connected and all integrable systems $F=(f_{1},\dots,f_{n}):M\to\mathbb{R}^{n}$ are proper with connected fibers.
A point $m\in M$ in which $X_{f_{1}}(m),...,X_{f_{n}}(m)$ of $T_{m}M$ are linearly dependent is said to be singular or a singularity of $F$. Otherwise we call it regular. The rank of $m$ is given by $\dim\operatorname{span}_{\mathbb{R}}\{X_{f_{1}}(m),...,X_{f_{n}}(m)\}$. If a fiber $F^{-1}(c)$ for some $c\in{\mathbb{R}}^{m}$ contains a singularity we call it singular and otherwise regular. Denote a $k$-dimensional torus briefly by $\mathbb{T}^{k}$. By the Liouville-Arnold-Mineur Theorem, see Arnold [arnol2013mathematical], a pre-image of a regular value is diffeomorphic to $\mathbb{T}^{n}$. Moreover, $\omega$ vanishes along regular fibers $F^{-1}(c)$, hence $F^{-1}(c)$ is a Lagrangian submanifold, traditionally called a Liouville torus. One often refers to an integrable system $(M,\omega,F)$ as a singular Lagrangian fibration where singular emphasizes that $F$ may have singular fibers of various kinds: for example tori of dimension $m\in\{0,...,n-1\}$, pinched-tori etc.
2.3. Action-angle coordinates
One important theorem about the symplectic geometry of integrable systems is the existence of so-called action-angle coordinates, see Arnold [arnol1963theorem]: each regular fiber, in addition to being diffeomorphic to $\mathbb{T}^{n}$, can be seen as sitting inside of the cotangent bundle $T^{*}\mathbb{T}^{n}$ as the zero section and, in a neighborhood of the fiber, the integrable system can be written in a symplectic normal form $F:T^{*}\mathbb{T}^{n}\rightarrow\mathbb{R}^{n}$:
Theorem 2.1.
(Liouville-Arnold-Mineur Theorem) Let $(M,\omega,F)$ be an integrable system and let $c\in F(M)$ be a regular value. If $\Lambda_{c}:=F^{-1}(c)$ is a regular, compact and connected fiber, then there exist neighborhoods $U\subset F(M)$ of $c$ and $V\subset\mathbb{R}^{n}$ of the origin, such that for
$\mathcal{U}:=\coprod_{r\in U}F^{-1}(r)\quad\text{and}\quad\mathcal{V}:=\mathbb{T}^{n}\times V\subset T^{*}\mathbb{T}^{n}$
we have that $(\mathcal{U},\omega|_{\mathcal{U}},F|_{\mathcal{U}})$ and $(\mathcal{V},\omega_{0}|_{\mathcal{V}},F|_{V})$ are isomorphic integrable systems, where $\omega_{0}$ is the standard symplectic form on $T^{*}\mathbb{T}^{n}$.
2.4. Linearization of non-degenerate singularities
A singular point $p\in M$ of rank zero of an integrable system $(M,\omega,F=(f_{1},...,f_{n}))$ is non-degenerate if the Hessians $d^{2}f_{1}(p),...,d^{2}f_{n}(p)$ span a so called Cartan subalgebra of the Lie algebra of quadratic forms on $T_{p}M$. We refer to Bolsinov & Fomenko [bolsinov2004integrable] for the precise definition of non-degenerate points of higher rank.
Non-degenerate singularities are in fact linearizable, more precisely, there exists a local normal form which is based on the works of Rüssmann [russmann1964verhalten], Vey [vey1978certains], Colin de Verdiére &\&& Vey [de1979lemme], Eliasson [eliasson1990normal, eliassonphdthesis], Dufour and Molino [dufour1991compactification], Miranda [miranda2014integrable, mirandaphdthesis], Miranda &\&& Zung [miranda2004equivariant], Miranda &\&& Vũ Ngǫc [miranda2005singular], Vũ Ngǫc &\&& Wacheux [vu2013smooth], and Chaperon [chaperon2013normalisation]: there are symplectic local coordinates $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ near the singular point $m$ in which $m=(0,...,0)$ corresponds to the origin, the symplectic form assumes the expression $\omega=\sum_{i=1}^{n}dx_{i}\wedge d\xi_{i}$ and there exist functions $q_{1},...,q_{n}$ of $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ such that the integrable system $F=(f_{1},...,f_{n})$ satisfies $\{f_{i},q_{j}\}=0$ for all indices $1\leq i,j\leq n$, where $q_{i}$ is one of the following functions:
  • elliptic type: $q_{i}=\frac{x_{i}^{2}+\xi_{i}^{2}}{2}$,
  • hyperbolic type: $q_{i}:=x_{i}\xi_{1}$,
  • real type: $q_{i}=\xi_{i}$,
  • focus-focus type: $q_{i}=x_{i}\xi_{i+1}-x_{i+1}\xi_{i}$ and $q_{i+1}=x_{i}\xi_{1}+x_{i+1}\xi_{i+1}$.
If there are no components of hyperbolic type then
$(F-F(m))\circ\phi=g\circ(q_{1},...,q_{n})$
where $\phi=(x_{1},...,x_{n},\xi_{1},...,\xi_{n})^{-1}$ and $g$ is a diffeomorphism from a small neighborhood of $(0,...,0)$ into another such neighborhood such that $g(0,...,0)=(0,...,0)$. In this paper, we will mostly work with singularities without components of hyperbolic type.
Throughout this paper, we always assume that the singularities of $F:M\rightarrow\mathbb{R}^{n}$ are non-degenerate.
2.5. Toric systems
Very ‘easy’ integrable systems are the following ones.
Definition 2.2.
An integrable system $(M,\omega,F=(f_{1},...,f_{n}))$ on a symplectic $2n$-dimensional manifold $(M,\omega)$ is toric if the Hamiltonian vector fields $X_{f_{1}},...,X_{f_{n}}$ generate periodic flows of the same period (in our convention $2\pi$) and the action of $\mathbb{T}^{n}$ on $M$ induced by these flows is effective (i.e. only the identity acts trivially).
Unless otherwise stated, we will usually consider toric integrable systems on closed connected manifolds.
The periodicity condition implies that the singularities of toric integrable systems cannot have focus-focus or hyperbolic components, that is, if $m=(0,...,0)$ and $\omega=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$, then the integrable system is locally in a neighborhood of $m$ of the form
$F(x_{1},...,x_{n},\xi_{1},...,\xi_{n})=\left(\frac{x_{1}^{2}+\xi_{1}^{2}}{2},...,\frac{x_{k}^{2}+\xi_{k}^{2}}{2},\xi_{k+1},...,\xi_{n}\right)$
for some $0\leq k\leq n$. Toric integrable systems have connected fibers, see Atiyah [atiyah1982convexity]. In particular, all fibers of a toric system $F$ are diffeomorphic to tori of varying dimensions $\mathbb{T}^{k}$ with $k\in\{0,...,n\}$. Note that this is not the case for general integrable systems.
One of the fundamental theorems of equivariant symplectic geometry, due to Atiyah [atiyah1982convexity] and Guillemin & Sternberg [guillemin1982convexity] says that, if $M$ is compact and connected, the image $F(M)$ is a convex polytope in $\mathbb{R}^{n}$, obtained as the convex hull of the images of the fixed points of the Hamiltonian action of the $n$-torus on $M$. This polytope has the property that if two toric integrable systems are isomorphic then their associated images coincide, possibly up to translations and composition with a matrix in $\text{GL}(n,\mathbb{Z})$. Furthermore, Delzant showed that the convex polytopes obtained as images of toric integrable systems are of a special type: they are simple, rational, and smooth. Such polytopes are called Delzant polytopes and being simple, rational, and smooth means that there are precisely $n$ edges meeting at each vertex, that the slopes of the edges are rational, and that the normal vectors to the facets meeting at each vertex form a basis of the integral lattice. Moreover, Delzant proved that such polytopes are in bijective correspondence with toric integrable systems on compact manifolds in the following sense:
  • Uniqueness: two systems $(M,\omega,F)$ and $(M^{\prime},\omega^{\prime},F^{\prime})$ are isomorphic if and only if they have the same convex polytope as image of the momentum map (up to translations and $\text{GL}(n,\mathbb{Z})$ transformations).
  • Existence: from any Delzant polytope $\Delta$ in $\mathbb{R}^{n}$ one can construct a toric integrable system $(M,\omega,F)$ on a compact connected symplectic $2n$-dimensional manifold with $F(M)=\Delta$. This construction uses the method of symplectic reduction recalled in Section 2.10.
Since $\Delta=F(M)$ classifies $(M,\omega,F)$, one can learn everything about $F$, up to isomorphisms, from $\Delta$. In particular the fiber structure of $F$ can be read off from the polytope $\Delta$: the fiber of $F$ over $p\in\Delta$ is diffeomorphic to a $k$-dimensional torus, where $k$ is the dimension of lowest dimensional face of $\Delta$ such that $p\in\Delta$.
2.6. Almost toric manifolds
So called almost toric manifolds, as introduced by Symington [symington2002four], are a generalization of closed toric manifolds and will be defined in more detail below.
Definition 2.3.
A locally trivial fibration of a $2n$-dimensional symplectic manifold $(M,\omega)$ is called a regular Lagrangian fibration if the fibers are Lagrangian submanifolds. More generally, a projection $\pi:(M,\omega)\rightarrow B$, where $B$ is an $n$-dimensional manifold with boundary, is a Lagrangian fibration if it restricts to a regular Lagrangian fibration over the open dense set $B_{r}\subset B$, defined as the set of regular values of $\pi$.
Definition 2.4.
By a lattice in a vector bundle of rank $n$ we mean a smooth varying lattice isomorphic to $\mathbb{Z}^{n}$ in each fiber of the bundle. An integral affine structure $\mathcal{A}$ on a $n$-manifold $B$ is a lattice in the tangent bundle of $B$. A manifold admitting such a structure is called an integral affine manifold.
The integral affine structure on the base of a regular Lagrangian fibration arises from a natural action of the cotangent bundle of the base on the total space: any $\alpha\in T^{*}B$ defines a vertical vector field $X_{\alpha}$ symplectically dual to $\alpha$, so for $x\in M$ the action of $\alpha\in T^{*}B$ is given by $\alpha\cdot x=\phi_{\alpha}(x)$, where $\phi_{\alpha}$ is the time-1 map of $X_{\alpha}$. The elements of the cotangent bundle of $B$ that act trivially, i.e., the isotropy group of this action, form a lattice $\Lambda^{*}$. The dual lattice $\Lambda$ in the tangent bundle defines the integral affine structure. We denote by $(\mathbb{R}^{n},\mathcal{A}_{0})$ the standard integral affine structure of $\mathbb{R}^{n}$ induced by $\mathbb{Z}^{n}$.
Definition 2.5.
Let $B$ be an integral affine manifold with integral affine structure $\mathcal{A}$ and $(M,\omega)$ a $2n$-dimensional symplectic manifold. A Lagrangian fibration $\pi:(M,\omega)\rightarrow B$ is a toric fibration if there exists a Hamiltonian $n$-torus action and an immersion $\Phi:(B,\mathcal{A})\rightarrow(\mathbb{R}^{n},\mathcal{A}_{0})$ such that $\Phi\circ\pi$ is the corresponding momentum map. Denote by $\mathcal{S}$ the following stratification: the $l$-stratum of $(B,\mathcal{S})$ is the set of points $b\in B$ such that $\pi^{-1}(b)$ is a torus of dimension $l$. We call the triple $(B,\mathcal{A},\mathcal{S})$ a toric base. The reduced boundary of a toric base $(B,\mathcal{A},\mathcal{S})$, denoted by $\partial_{R}B$, is the set of points in $B$ that belong to the lower-dimensional strata, i.e., $k$-strata for $k<n$.
From a toric base $(B,\mathcal{A},\mathcal{S})$ it is possible to reconstruct a toric fibration and a symplectic manifold by starting with a regular Lagrangian fibration over the base $(B,\mathcal{A})$ and then collapse certain fibers to lower dimensional tori so as to obtain the stratification $\mathcal{S}$. The collapsing of the fibers is achieved via boundary reduction, see Symington [symington2002four, Proposition $3.8$] for more details.
Definition 2.6.
Let $\pi:(M,\omega)\rightarrow(B,\mathcal{A},\mathcal{S})$ be a toric fibration and $\gamma$ the image in $B$ of a compact embedded curve with one endpoint $b_{1}$ in the $(n-1)$-dimensional stratum of $\partial_{R}B$ and such that $\gamma\setminus\{b_{1}\}\subset B_{r}$. Let $b_{2}\in B$ be the other endpoint of $\gamma$. The collapsing class with respect to $\gamma$ for the smooth component of $\partial_{R}B$ containing $b_{1}$ is defined as the class $a\in H_{1}(F_{b_{2}},\mathbb{Z})$ that spans the kernel of $i_{*}:H_{1}(F_{b_{2}},\mathbb{Z})\rightarrow H_{1}(\pi^{-1}(\gamma),\mathbb{Z})$ where $i$ is the inclusion map $i:F_{b_{2}}\rightarrow\pi^{-1}(\gamma)$. Accordingly, the collapsing covector with respect to $\gamma$ is the covector $v^{*}\in T^{*}_{b_{2}}B$ that determines vectors $v(x)\in T^{vert}_{x}M:=\ker(d_{x}\pi)$ for each $x\in\pi^{-1}(B_{r})$ such that the flow lines of this vector field represent $a$. (Notice $a$ and $v^{*}$ are well-defined up to sign and have coefficients in $\mathbb{Z}$.)
Remark 2.7.
If $v$ is a primitive integral vector that is normal with respect to the standard Euclidean metric to a smooth component of the reduced boundary of $(B,\mathcal{A}_{0},\mathcal{S})\subset(\mathbb{R}^{2},\mathcal{A}_{0})$ then $v^{*}$ is a collapsing covector for that component.
Definition 2.8.
An almost toric fibration of a symplectic $2n$-manifold $(M,\omega)$ is a Lagrangian fibration $\pi:(M,\omega)\rightarrow B$ such that any critical point of $\pi$ has a Darboux neighborhood (i.e. local coordinates as in Section 2.4) in which the components of the projection $\pi=(\pi_{1},...,\pi_{k},\pi_{k+1},...,\pi_{n})$ are of the form $\pi_{j}(x,y)=x_{j}$ for $j\leq k$ for some $k\leq 2n$ and otherwise are of the following two forms:
  • elliptic (or toric) : $\pi_{j}(x,y)=(x_{j}^{2}+y_{j}^{2})$,
  • nodal (or focus-focus): $(\pi_{i},\pi_{i+1})(x,y)=(x_{i}y_{i+1}-x_{i+1}y_{i},\ x_{i}y_{i}+x_{j}y_{j})$.
An almost toric manifold is a symplectic manifold equipped with an almost toric fibration.
Assigning points in the base of an almost toric fibration to strata according to the dimension of their preimage yields a stratification $\mathcal{S}$ of the base. When $(M,\omega)$ has dimension $4$ the images of nodal singular points (nodes) are isolated points that belong to the top dimensional stratum. Let $\Sigma$ be the codimension two set of points containing a nodal singularity in their preimage.
By the integral affine structure $\mathcal{A}$ on the base of an almost toric fibration we mean the affine structure defined on $B\setminus\Sigma$. If an affine structure $\mathcal{A}$ and stratification $\mathcal{S}$ are induced from an almost toric fibration, we call the triple $(B,\mathcal{A},\mathcal{S})$ an almost toric base.
Let $\text{Aff}(n,\mathbb{Z})$ be the group of integral affine transformations, i.e., maps $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ of the form $T(x)=Ax+b$, with $A\in GL(n,\mathbb{Z})$ and $b\in\mathbb{R}^{n}$.
The affine monodromy of the associated integral affine manifold $B_{r}$ is defined as follows. Let $\Lambda$ be the lattice in $TB_{r}$. Choose a point $b\in B_{r}$, identify $(T_{b}B_{r},\Lambda_{b})$ with $(\mathbb{R}^{n},\mathbb{Z}^{n})$ and for each element $\alpha\in\pi_{1}(b,B_{r})$ choose a representative $\gamma_{\alpha}:I:=[0,1]\rightarrow B_{r}$. The monodromy representation relative to these choices is denoted by $\Psi_{B}:\pi_{1}(b,B_{r})\rightarrow\text{Aff}(n,\mathbb{Z})$ and defined as follows: for all $\alpha\in\pi_{1}(b,B_{r})$, one defines $\Psi_{B_{r}}(\alpha)$ as the automorphism of $(\mathbb{R}^{n},\mathbb{Z}^{n})$ such that $\gamma_{\alpha}^{*}(TB_{r},\Lambda)$ is isomorphic to $I\times(\mathbb{R}^{n},\mathbb{Z}^{n})/\left((0,p)\simeq(1,\Psi_{B_{r}}(\alpha)(p))\right)$ for $p\in\mathbb{R}^{n}$. The monodromy is the equivalence class of monodromy representations relative to different points in $B$ and choices of identifications of $(T_{b}{B_{r}},\Lambda_{b})$ with $(\mathbb{R}^{n},\mathbb{Z}^{n})$. Analogously, for a almost toric fibration $\pi:(M,\omega)\rightarrow B$ one can define the topological monodromy of $\alpha\in\pi_{1}(b,B_{r})$ by considering $\pi_{1}(\pi^{-1}(b),\mathbb{Z})\cong\mathbb{Z}^{n}$ instead of $\Lambda_{b}$.
Definition 2.9.
Let $\pi:(M,\omega)\rightarrow B$ be an almost toric fibration that has a node at $s\in B$. Let $\eta$ be the image in $B$ of an embedded curve with endpoints at $s$ and at a regular point $b\in B_{r}$ such that $\eta\setminus\{s\}\subset B_{r}$ contains no other nodes. Associated to $s$ and $\eta$ is the (well defined up to sign) vanishing class in $H_{1}(F_{b},\mathbb{Z})$, namely the class whose representatives bound a disk in $\pi^{-1}(\eta)$. The vanishing covector $v^{*}\in T_{b}^{*}B$ is the covector (defined up to sign) that determines vectors $v(x)\in T_{x}^{vert}M$ for each $x\in\pi^{-1}(b)$ such that the flow lines of this vector field represent the vanishing class.
Lemma 2.10.
(Symington, [symington2002four, Lemma 4.10]) Using the notation from Definition 2.9 suppose $\gamma:S^{1}\rightarrow B$ is a positively oriented loop based at $b\in B$ that is the boundary of a closed neighborhood of $s\in B$ containing $\eta\subset B$. Then the vanishing class, up to rescaling, is the unique class that is preserved by the monodromy along $\gamma$.
Let $\pi:(U,\omega)\rightarrow B$ be a Lagrangian fibered neighborhood of a nodal fiber in a $4$-manifold where $B$ is connected and (the value of) the node is $s\in B$. Let $\mathcal{A}$ be the affine structure on $B\setminus\{s\}$. The affine monodromy is non-trivial, due to the presence of the node (Symington, [symington2002four] Section $4.3$). Hence there is no immersion of $(B\setminus\{s\},\mathcal{A})$ into $(\mathbb{R}^{2},\mathcal{A}_{0})$. However, there is an affine immersion if one removes a curve from $B^{2}$ that connects the (value of the) node to the boundary since the remaining set is simply connected. We define the base diagram for the neighborhood of a node, as the image of an affine immersion of the complement of such a curve.
Definition 2.11.
Let $\pi:(M,\omega)\rightarrow B$ be a an almost toric fibration. A $1$-dimensional submanifold is affine linear if at every point it has a tangent vector in $\Lambda$. The eigenline $L\subset B$ through a node $s$ is the unique maximal affine linear immersed one-manifold through the node for which there is a regular point $b\in L$, arbitrarily close to $s$, such that the affine monodromy along an arbitrarily small loop around $s$ and based at $b$ preserves $T_{b}L\subset T_{b}B$. An eigenray is either of the two maximal affine linear submanifolds that has an endpoint at the node and is a subset of the eigenline.
Remark 2.12.
Consider a base diagram for an almost toric base $(B,\mathcal{A},\mathcal{S})$ of an almost toric fibration. If a primitive integral vector $v$ is perpendicular to the eigenline through a node then $v^{*}$ is a covector that defines the vanishing class of the fiber, with respect to the eigenline.
2.7. Nodal trades and slides
Henceforth we work with $4$-dimensional symplectic manifolds $(M,\omega)$, therefore appearing bases $B$ are $2$-dimensional in this subsection. Nodal trading and sliding are surgery operations that change an almost toric fibration of a symplectic $4$-manifold into another almost toric fibration of the same symplectic manifold. In a nodal trade the essential idea is that one can trade an elliptic-elliptic fiber for a nodal fiber, and vice versa under the appropriate conditions. A nodal slide should be thought of as a one-parameter family of almost toric bases in which a node moves in the base along its eigenline.
Definition 2.13.
Two almost toric bases $(B,\mathcal{A}_{i},\mathcal{S}_{i}),i=1,2$, are related by a nodal slide if there is a curve with image $\gamma\subset B$ such that $(B\setminus\gamma,\mathcal{A}_{1},\mathcal{S}_{1})$ and $(B\setminus\gamma,\mathcal{A}_{2},\mathcal{S}_{2})$ are isomorphic (i.e. there exists a map $\phi:(B\setminus\gamma,\mathcal{A}_{1},\mathcal{S}_{1})\rightarrow(B\setminus\gamma,\mathcal{A}_{2},\mathcal{S}_{2})$ that preserves the affine structures and stratifications), and for each $i$, the curve $\gamma$ contains one node of $(B,\mathcal{A}_{i},\mathcal{S}_{i})$ and $\gamma$ belongs to the eigenline through that node.
Proposition 2.14.
(Symington, [symington2002four, Proposition 6.2]) If two bases are related by a nodal slide then they define the same manifold equipped with isotopic symplectic structures. Furthermore, the manifolds are fiberwise symplectomorphic on the complement of a compact set belonging to the preimage of the eigenline.
Lemma 2.15.
(Symington, [symington2002four, Lemma 6.3]) Let $(B,\mathcal{A},\mathcal{S})$ be an almost toric base and $R\subset B$ be the image of an embedded eigenray that connects a node $s\in B$ with a point $b$ in an edge $E\subset\partial_{R}B$ such that there are no other nodes on $R$. Let $v^{*}$, $w^{*}\in T_{b}(B)^{*}$ be the vanishing and collapsing covectors. If $v^{*}$ and $w^{*}$ span $\Lambda_{b}^{*}$ then there is an almost toric base $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ such that:
  • $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ contains one node less than $(B,\mathcal{A},\mathcal{S})$,
  • $(B\setminus R,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ is isomorphic to $(B\setminus R,\mathcal{A},\mathcal{S})$,
  • in $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$, the intersection of $R$ and $\partial_{R}B$ is a vertex.
In this situation we say that the two almost toric basis $(B,\mathcal{A},\mathcal{S})$ and $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ are related by a nodal trade.
Theorem 2.16.
(Symington [symington2002four, Theorem 6.5]) Suppose two almost toric bases that are related by a nodal trade. Then the symplectic manifolds induced by the almost toric bases are symplectomorphic. In fact, their symplectic structures are isotopic.
2.8. Semitoric systems
The class of toric systems can be generalized to semitoric systems, which are a special case of an almost toric fibration.
Definition 2.17.
An integrable system
$(M,\omega,F=(f_{1},...,f_{n}):M\rightarrow\mathbb{R}^{n})$
is semitoric if the Hamiltonian vectors fields $X_{f_{1}},...,X_{f_{n-1}}$ generate periodic flows of the same period (in our convention $2\pi$), if the action of $\mathbb{T}^{n-1}$ induced on $M$ by these flows is effective and if the singularities of $F$ are non-degenerate and do not have hyperbolic components. If $M$ is not compact the integrals $f_{1},...,f_{n-1}$ are required to be proper.
The singular fibers of these systems are either points, circles or tori with a finite number of pinches. Note that this last type of fiber does not appear in toric systems.
If $M$ is $4$-dimensional, semitoric systems were classified, first under some conditions, by Pelayo &\&& Vũ Ngǫc [pelayo2009semitoric, pelayo2011constructing] and then in full generality by Palmer & Pelayo & Xiudi [palmer2019semitoric]. The distinguishing property of a semitoric system is that all but maybe one of the integrals generate periodic flows. So toric systems are a special case of semitoric ones. We will consider examples in Sections 4.1 & 4.2 & 4.3.
A simple semitoric system is a semitoric system where each fiber contains at most one focus-focus singularity. Such systems are determined up to isomorphism, according to Pelayo &\&& Vũ Ngǫc [pelayo2009semitoric], by the following $5$ symplectic invariants:
  1. (1)
    The number of focus-focus singularities, denoted by $n_{FF}$.
  2. (2)
    The Taylor series invariant, $n_{FF}$ formal Taylor series in two variables describing the foliation around each focus-focus singular fiber.
  3. (3)
    The polytope invariant, a family of weighted rational convex polytopes (generalizing the Delzant polytope of toric systems).
  4. (4)
    The height invariant, given by $n_{FF}$ numbers corresponding to the height of the focus-focus critical values in the rational convex polytopes of the polytope invariant.
  5. (5)
    The twisting index invariant, given by $n_{FF}$ integers measuring how twisted the system is around singularities from a ‘toric point of view’.
Let $(M,\omega,(f_{1},f_{2}))$ and $(\tilde{M},\tilde{\omega},(\tilde{f}_{1},\tilde{f}_{2}))$ be two semitoric systems. We say that they are isomorphic as semitoric systems if there exists a symplectomorphism $\psi:M\rightarrow\tilde{M}$ and a smooth map $g:\mathbb{R}^{2}\rightarrow\mathbb{R}$ such that $(\tilde{f}_{1},\tilde{f}_{2})\circ\psi=(f_{1},g(f_{1},f_{2}))$ and $\partial_{2}g>0$.
The classification result for semitoric systems has two aspects:
  • Two semitoric systems are isomorphic if and only if they have the same list of symplectic invariants
  • Given any admissible list of invariants, a semitoric system with such invariants can be constructed.
Since we will later work with (representatives of) the polytope invariant, let us consider it now in more detail. Let
$\{c_{i}=(x_{i},y_{i})\mid i=1,...,m_{f}\}\in\mathbb{R}^{2}$
[…]
$l^{\vec{\epsilon}}=\cup_{i}l_{i}$
where in addition we decorate each $l_{i}$ with the multiplicity $\epsilon_{i}k_{i}$, where $k_{i}$ is the number of focus-focus points in the fiber $F^{-1}(c_{i})$. If several $c_{i}$ have the same $x_{i}$ coordinate, then $l_{i}$ is the union of all corresponding segments. Given $c\in l_{i}$, define $k(c):=\sum_{c_{j}}\epsilon_{j}k_{j}$ where the sum runs over all focus-focus values $c_{j}$ such that $c_{i}\in l_{j}$.
Denote by $\mathcal{T}$ the subgroup of $\text{Aff}(2,\mathbb{Z})$ which leaves a vertical line, with orientation, invariant. In other words an element of $\mathcal{T}$ is a composition of a vertical translation and an element of $\{T^{k},k\in\mathbb{Z}\}\subset GL(2,\mathbb{Z})$, where
$T^{k}:=\begin{bmatrix}1&0\\ k&1\end{bmatrix}.$
Theorem 2.18.
(Vũ Ngǫc, [vu2007moment, Theorem 3.8]) Using the notation from above for a semitoric system $(M,\omega,(f_{1},f_{2}))$, for all $\vec{\epsilon}\in\{-1,+1\}^{m_{f}}$, there exists a homeomorphism $f_{\vec{\epsilon}}=(f_{\vec{\epsilon}}^{(1)},f_{\vec{\epsilon}}^{(2)})$ from $B$ to $f_{\vec{\epsilon}}(B)\subseteq{\mathbb{R}}^{2}$ such that
  • $f_{\vec{\epsilon}}|_{(B\setminus l^{\vec{\epsilon}})}$ is a diffeomorphism onto its image;
  • $f_{\vec{\epsilon}}|_{(B\setminus l^{\vec{\epsilon}})}$ is affine;
  • $f_{\vec{\epsilon}}$ preserves $f_{1}$, i.e., $f_{\vec{\epsilon}}(x,y)=(x,f_{\vec{\epsilon}}^{(2)}(x,y))$;
  • For all $i\in\{1,...,m_{f}\}$ and all $c\in\text{Int}(l_{i})$, $f_{\vec{\epsilon}}|_{(B\setminus l^{\vec{\epsilon}})}$ extends to a smooth map in the domain $\{(x,y)\in D|\ x\leq x_{i}\}$ and $\{(x,y)\in D|\ x\geq x_{i}\}$, where $D$ is an open ball around $c$. Furthermore,
    $\lim_{\stackrel{{\scriptstyle(x,y)\rightarrow c}}{{x<x_{i}}}}df_{\vec{\epsilon}}(x,y)=T^{k(c)}\lim_{\stackrel{{\scriptstyle(x,y)\rightarrow c}}{{x>x_{i}}}}df_{\vec{\epsilon}}(x,y)$
where
$T^{k(c)}=\begin{bmatrix}1&0\\ k(c)&1\end{bmatrix}.$
Such an $f_{\vec{\epsilon}}$ is unique modulo a left composition by a transformation in $\mathcal{T}$, and its image is a representative of the polytope invariant.
The map $f_{\vec{\epsilon}}$ is sometimes referred to as straightening homeomorphism. In order to arrive at $f(B)$, one cuts the set $B$ along each of the vertical lines $\mathcal{L}_{i}^{\epsilon_{i}}$ to the focus-focus values. Then the resulting image becomes simply connected, and thus there exists a global $2$-torus action on the preimage of this set. A representative of the polytope invariant can thus be seen as closure of the image of a toric momentum map.
Describing all possible choices (like for instance the signs $\epsilon_{j}$) by means of a group action allows to write the polytope invariant as equivalence class or orbit of this group action. Since we work later always with representatives, we refer for details to Vũ Ngǫc [vu2007moment].
2.9. Semitoric systems and almost toric fibrations
Let $(M,\omega,F)$ be a semitoric system. In particular, it induces an almost toric fibration and an almost toric base. Let $\Delta_{\vec{\epsilon}}$ be a representative of the polytope invariant for this semitoric system. Then, by Vũ Ngǫc [vu2007moment], the eigenline of a nodal value in $\Delta_{\vec{\epsilon}}$ is the vertical line that passes through the node. The affine monodromy can be represented, with respect to an integral affine basis of $\Delta_{\vec{\epsilon}}$, as
$\begin{bmatrix}1&0\\ k&1\end{bmatrix}$
for some $k\in\mathbb{N}$, where $k$ is the number of focus-focus critical points in the critical fiber.
Definition 2.19.
Let $p$ be a corner of a polytope in $\mathbb{R}^{n}$. We say that $p$ is a Delzant corner if there exists an integral affine transformation $A$ that maps $p$ and its adjacent edges to the corner at $(0,0)$ with adjacent edges created by $(1,0)$ and $(0,1)$, which is called the standard Delzant corner. This is equivalent to the fact that there exist vectors spanning the edges of $p$ that are a $\mathbb{Z}$ basis of $\mathbb{Z}^{n}$.
Remark 2.20.
Let $\Delta_{\vec{\epsilon}}$ be a representative of the polytope invariant for a simple semitoric system $(M,\omega,F)$ and $R$ be an eigenray of $\Delta_{\vec{\epsilon}}$ that connects the node to a point $p$ in the boundary which is a Delzant corner. Then Lemma 2.15 applies, and we are able to do a nodal trade.
2.10. Symplectic reduction
Let $(M,\omega)$ be a symplectic toric manifold of dimension $2N$ with $\mathbb{T}^{N}$-action generated by a momentum map $\mu:M\rightarrow(\mathbb{R}^{N})^{*}$ and set $P:=\mu(M)$.
In this section we recall the following symplectic reduction result from Abreu & Macarini [abreu2013remarks]: let $x\in\text{int}(P)$ and let $T_{x}:=\mu^{-1}(x)$ denote the corresponding $\mathbb{T}^{N}$ orbit (which is a Lagrangian torus in $M$). Let $K\subset\mathbb{T}^{N}$ be a subtorus of dimension $N-n\geq 0$ determined by an inclusion of Lie algebras $\imath:\mathbb{R}^{N-n}\rightarrow\mathbb{R}^{N}$. The momentum map for the induced action of $K$ on $M$ is given by
$\mu_{K}=\imath^{*}\circ\mu:M\rightarrow(\mathbb{R}^{N-n})^{*}.$
Let $c\in\mu_{K}(M)\subset(\mathbb{R}^{N-n})^{*}$ be a regular value and assume that $K$ acts freely on the level set $Z:=\mu_{K}^{-1}(c)\subset M$. Then, the reduced space $(M^{\text{red},c}:=Z/K,\omega^{\text{red},c})$ is a symplectic toric manifold of dimension $2n$ with $\mathbb{T}^{\text{red},c}:=\mathbb{T}/K$ action generated by the momentum map
$\mu^{\text{red},c}:M^{\text{red},c}\rightarrow P^{\text{red},c}\subset(\mathbb{R}^{n})^{*}\cong\ker(\imath^{*}).$
The symplectic form $\omega^{\text{red},c}$ is characterized by
$\pi^{*}(\omega^{\text{red},c})=\omega|_{Z}$
where $\pi:Z\rightarrow Z/K$.
2.11. Displacing toric fibers
In McDuff [mcduff2011displacing] conditions are given to displace Lagrangian toric fibers of toric manifolds. In this section, we summarize the necessary notions and results of McDuff [mcduff2011displacing] which we will need later on.
The affine distance $d_{\text{aff}}(x,y)$ between two points $x$, $y$ on a line $L$ with rational slope is defined as the quotient of the Euclidean distance $d_{E}(x,y)$ and the minimum Euclidean distance from 00 to an integral point $p\in\mathbb{Z}^{n}$ on the line through 00 parallel to $L$. Equivalently, if $\phi$ is any integral affine transformation of $\mathbb{R}^{n}$ that takes $x,y$ to the first coordinate axis (the $x_{1}$-axis) then $d_{\text{aff}}(x,y)=d_{E}(\phi(x),\phi(y))$. An affine line $L=z+\mathbb{R}v$ in $\mathbb{R}^{n}$ is rational if the direction vector $v$ can be taken to be a primitive integral vector in $\mathbb{Z}^{n}$. Given an affine rational line $L$ with primitive direction $v\in\mathbb{Z}^{n}$, the affine distance $d_{\text{aff}}$ between two points $x,y\in L$ is $t\in\mathbb{R}$ such that $x-y=tv\in\mathbb{R}^{n}$. An affine hyperplane $A$ is called rational if it has a primitive integral normal vector $\eta$, i.e., if it is given by an equation of the form $\langle x,\eta\rangle=k$ where $k\in\mathbb{R}$ and $\eta$ is primitive and integral. The affine distance $d_{\lambda}(x,A)$ from a point $x$ to a rational affine hyperplane in the rational direction $\lambda$ is defined as
(2.1) $d_{\lambda}(x,y):=d_{\text{aff}}(x,y)$
where $y\in A$ lies on the ray $x+a\lambda$ with $a\in\mathbb{R}^{+}$. If the ray does not meet $A$ we set $d_{\lambda}(x,A)=+\infty$. We say that an integral vector $\lambda$ is integrally transverse to $A$ if $|\langle\lambda,\eta\rangle|=1$ where $\eta$ is the normal as above.
Definition 2.21.
Let $w$ be a point on a facet $F$, i.e., on a $(n-1)$-dimensional face of a rational polytope $\Delta$ and let $\lambda\in\mathbb{Z}^{n}$ be integrally transverse to $F$. The probe $p_{F,\lambda}(w)=:p_{\lambda}(w)$ with direction $\lambda\in\mathbb{Z}^{n}$ and initial point $w\in F$ is the half open line segment consisting of $w$ together with the points in $\text{int}(\Delta)$ that lie on the ray from $w$ in direction $\lambda$.
In the next lemma we could use any notion of length, but the affine distance is the most natural one.
Lemma 2.22.
(McDuff [mcduff2011displacing, Lemma 2.4]) Let $\Delta$ be a smooth momentum polytope associated with a symplectic toric manifold $(M,\omega,F)$, i.e., $\Delta=F(M)$. Let $u\in\text{int}(\Delta)$ lie on the probe $p_{F,\lambda}(w)$. If $w$ lies in the interior of its corresponding facet and if $u$ is less than halfway measured from the boundary along $p_{\lambda}(w)$, then the fiber $F^{-1}(u)$ is displaceable.
Definition 2.23.
Let $\Delta$ be any rational polytope and $u\in\text{int}(\Delta)$. If there is a probe $p_{F,\lambda}(w)$ through $u$ that satisfies the conditions of Lemma 2.22 then $u$ is said to be displaceable by the probe $p_{F,\lambda}(w)$.
2.12. Special Hamiltonian diffeomorphisms of the disk
In this subsection, we recall a result from Abreu &\&& Borman &\&& McDuff [abreu2014displacing] that concerns Hamiltonian diffeomorphisms of the disk which we will need to displace focus-focus fibers in Section 3.3.
For real numbers $0<A<B$, consider a smooth, non-decreasing function $a:[0,B-A]\rightarrow[A,B]$ such that the function $s\mapsto a(s)+s$ is non-decreasing for $s\in[0,B-A]$ and
$A\leq a(s)+s\leq B\quad\mbox{and}\quad a(B-A)=A.$
For our purposes, we pick $a(s):=B-s$. For $r>0$, let $\mathbb{D}(r)\subset\mathbb{R}^{2}$ denote the $2$-dimensional disk of radius $r$ centered at the origin.
Lemma 2.24 (Abreu &\&& Borman &\&& McDuff, [abreu2014displacing, Lemma 6.4.2]).
For any function $a(s)$ as above and any $\epsilon>0$, there is a compactly supported Hamiltonian diffeomorphism $\rho:\text{Int}(\mathbb{D}(B))\rightarrow\text{Int}(\mathbb{D}(B))$ such that
$\rho(\mathbb{D}(s))\subset\mathbb{A}(a(s),a(s)+s+\epsilon),\qquad\forall\ \epsilon\leq s\leq B-A-\epsilon,$
where $\mathbb{A}(b,c):=\mathbb{D}(b)\backslash\text{Int}(\mathbb{D}(b))$.
Corollary 2.25.
If $A>\frac{B}{2}$, every circle of radius less than $B-A$ is displaced by $\rho$.
2.13. Symplectic quasi-states
Symplectic quasi-states are a useful tool to find nondisplaceable subsets in a symplectic manifold, and so are the notions of superheavy, heavy and pseudoheavy sets. Entov &\&& Polterovich [entov2009rigid, polterovich2006quasi] introduced the notions of heaviness and superheaviness. Furthermore, Kawasaki & Orita [kawasaki2021existence] added the notion of pseudoheaviness. The definition of a symplectic quasi-state and pseudoheavy, heavy, superheavy subsets is based on spectral invariants in Hamiltonian Floer theory (see Oh [oh2005construction]). We now recall in more detail some definitions and results from these papers that we need later on.
In what follows we require all symplectic manifolds $(M,\omega)$ to be closed. However we note that the theory of symplectic quasi-states could be developed for non-compact symplectic manifolds by considering compactly supported Hamiltonian functions.
Definition 2.26.
A partial symplectic quasi-state on a symplectic manifold $(M,\omega)$ is a map $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ that satisfies the following conditions:
  • Normalization: $\zeta(a)\equiv a$ for all $a\in\mathbb{R}$.
  • Stability: For all $H_{1},H_{2}\in C^{\infty}(M)$:
    $\min_{M}(H_{1}-H_{2})\leq\zeta(H_{1})-\zeta(H_{2})\leq\max_{M}(H_{1}-H_{2}).$
  • Semi-Homogenity: $\zeta(sH)=s\zeta(H)$ for all $H\in C^{\infty}(M)$ and all $s>0$;
  • Hamiltonian invariance: $\zeta(H\circ\phi)=\zeta(H)$ for all $H\in C^{\infty}(M)$ and all $\phi\in\text{Ham}(M,\omega)$.
  • Vanishing: $\zeta(H)=0$ for all $H\in C^{\infty}(M)$ with support displaceable from itself.
[…]
Definition 2.27.
Symplectic quasi-states are partial symplectic quasi-states that satisfy the following quasi-linearity assumption: if $\{H,K\}=0$ then $\zeta(H+aK)=\zeta(H)+a\zeta(K)$ for $a\in\mathbb{R}^{>0}$.
Definition 2.28.
A map $\zeta:C^{\infty}(M,\omega)\rightarrow\mathbb{R}$ satisfies the so-called PB-inequality if there exists $C>0$ such that
$|\zeta(H+K)-\zeta(H)-\zeta(K)|\leq C\sqrt{||\{H,K\}||}\quad\forall\ H,K\in C^{\infty}(M,\omega).$
Definition 2.29.
[…]
$\zeta(H)\geq\inf_{X}H\quad(resp.\hskip 5.69054pt\zeta(H)\leq\sup_{X}H)$
[…]
These sets have important properties and relations between them:
Theorem 2.30.
(Kawasaki &\&& Orita [kawasaki2021existence, Proposition 1.8], Entov &\&& Polterovich [entov2009rigid, Theorem $1.4$]) Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$:
  1. (1)
    Every $\zeta$-superheavy subset is $\zeta$-heavy.
  2. (2)
    Every $\zeta$-heavy subset is $\zeta$-pseudoheavy.
  3. (3)
    Every $\zeta$-pseudoheavy subset is nondisplaceable from itself.
  4. (4)
    Every $\zeta$-pseudoheavy subset is nondisplaceable from every $\zeta$-superheavy subset.
One important result is that there always exists a pseudoheavy fiber for a partial symplectic quasi-state on an integrable system $(M,\omega,F):$
Theorem 2.31.
(Kawasaki &\&& Orita, [kawasaki2021existence, Theorem 1.7]) Let $(M,\omega,F)$ be an integrable system. Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$. Then there exists $y_{0}\in F(M)$ such that $F^{-1}(y_{0})$ is $\zeta$-pseudoheavy.
Remark 2.32.
Entov & Polterovich [polterovich2006quasi] showed the existence of a partial symplectic quasi-state for $2n$-dimensional symplectic manifolds which are rational and strongly semi-positive. Due to the work of Oh [oh2009floer] and Usher [dualityusher, spectralusher] the conditions can be dropped, see Entov [entovclarification] for more details. Note that all symplectic 4-manifolds are strongly semi-positive.
The existence of partial symplectic quasi-states on symplectic 4-manifolds thus implies:
Theorem 2.33.
(Entov &\&& Polterovich, [polterovich2006quasi, Theorem 2.1], Usher [dualityusher, spectralusher], Oh [oh2009floer]) Any integrable system $(M,\omega,F)$ on a symplectic $4$-manifold $(M,\omega)$ has a nondisplaceable fiber.
Definition 2.34.
Let $(M,\omega,F)$ be an integrable system and $p\in F(M)$. A fiber $F^{-1}(p)$ is called a stem if all other fibers of $F$ are displaceable.
Theorem 2.35.
(Polterovich &\&& Rosen, [polterovich2014function, Corollary 6.1.6]) Any stem is nondisplaceable by the group of symplectomorphisms of $(M,\omega)$.
Borman [borman2013quasi] showed that symplectic quasi-states behave well under symplectic reduction on a superheavy subset: let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold equipped with a smooth map $\Phi=(\Phi_{1},\cdot\cdot\cdot,\Phi_{k}):W\rightarrow\mathbb{R}^{k}$ and a regular level set $Z:=\Phi^{-1}(0)$. Suppose that all $\Phi_{i}$ mutually Poisson commute at each point in $Z$ and that $\Phi$ induces a free Hamiltonian $\mathbb{T}^{k}$-action on $Z$. Let $(M^{\text{red},0}:=Z/\mathbb{T}^{k},\omega^{\text{red},0})$ be the reduced space when performing symplectic reduction at 00. Denote by $\pi:Z\rightarrow M^{\text{red},0}$ the quotient map.
Theorem 2.36.
(Borman, [borman2013quasi, Theorem 1.1]) Let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold and $Z$ has above. If $\zeta:C^{\infty}(W,\omega)\rightarrow\mathbb{R}$ is a symplectic quasi-state satisfying the PB-inequality and $Z$ is superheavy for $\zeta$, then $\zeta$ induces a symplectic quasi-state
$\zeta^{\text{red},0}:C^{\infty}(M^{\text{red},0},\omega^{\text{red},0})\rightarrow\mathbb{R}$
satisfying the PB-inequality. If $Y\subset Z$ is superheavy for $\zeta$ then $\pi(Y)$ is superheavy for $\zeta^{\text{red},0}.$
3. Displacing focus-focus fibers
3.1. Generalization of the method of probes
In this subsection, we generalize the method of probes, which we will apply in subsequent sections to displace focus-focus fibers.
Lemma 3.1.
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,h],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, and $R>0$ and $0<h<\frac{R}{2}$. Then the set $F^{-1}([0,h]\times[0,h])\subset M$ is displaceable.
Proof.
Let us denote the integral normal vectors by $v_{1}=(1,0)$ and $v_{2}=(0,1)$ and notice that $\langle v_{1},v_{1}\rangle=\langle v_{2},v_{2}\rangle=1$ and $\langle v_{1},v_{2}\rangle=0$. Since we are in a toric fibration we have well-defined action-angle coordinates $(x,\theta)\in B\times\mathbb{T}^{2}$. Using these action-angle coordinates and the assumptions in the statement, possibly after applying an integral affine transformation, we have an embedding
$\Psi:\mathbb{D}(R)\times\mathbb{D}(h)\rightarrow(M,\omega),\qquad(s,\theta_{1},t,\theta_{2})\mapsto(sv_{1}+tv_{2},\theta_{1}v_{1}+\theta_{2}v_{2})\in B\times\mathbb{T}^{2}$
where $s$ and $t$ parametrize the radii of the disks. Let $\pi_{i}$ be the projections onto the disks for $i=1,2$. Then $\Psi^{*}\omega=\pi_{1}^{*}(\omega_{0})+\pi_{2}^{*}(\omega_{0})$.
Since $h<\frac{R}{2}$ we can use the results of Section 2.12 to obtain a compactly supported Hamiltonian $G_{h}:\mathbb{D}(R)\rightarrow\mathbb{R}$ that displaces every circle of radius less than or equal to $h$. This Hamiltonian allows us to define the compactly supported Hamiltonian
$G:\mathbb{D}(R)\times\mathbb{D}(h)\quad\rightarrow\quad\mathbb{R},\qquad G:=G_{h}\circ\pi_{1}.$
Since this Hamiltonian does not depend on $\mathbb{D}(h)$ and $\Psi^{*}\omega=\pi_{1}^{*}(\omega_{0})+\pi_{2}^{*}(\omega_{0})$, the flow of $G\circ\Psi^{-1}$ displaces the set $F^{-1}([0,h]\times[0,h])$ inside $\text{Im}(\Psi)$. Since $H$ is compactly supported we can extend it to the whole manifold $(M,\omega)$. ∎
The above proof can be adapted to obtain the following corollary:
Corollary 3.2.
Let $F:(M,\omega)\rightarrow B$ be a toric fibration. Suppose that there exists an affine embedding of $([0,R]\times[0,a],\mathcal{A}_{0})$ into $(B,\mathcal{A})$ where $\mathcal{A}$ is the integral affine structure induced by $F$ on $B$, and $R,a>0$, $0<h<\frac{R}{2}$. Then the set $F^{-1}([0,h]\times[0,a])\subset M$ is displaceable.
3.2. Nondisplaceability of focus-focus fibers with higher multiplicity
Let $c$ be a focus-focus value of a semitoric system $F$ on a $4$-dimensional symplectic manifold $(M,\omega)$, and let $k$ be the number of focus-focus point on the fiber $F^{-1}(c)$, often called the multiplicity of the focus-focus fiber $F^{-1}(c)$. Intuitively, a fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of $k$ spheres joint at the poles, see Zung [zungarnold] and Vũ Ngǫc [vu2000bohr, Section $6$]. In particular, these are Lagrangian spheres. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded.
Proposition 3.3.
An embedded Lagrangian sphere in a symplectic $4$-manifold is nondisplaceable.
Proof.
Recall the notion of intersection number and Euler characteristic (for a reference see Hirsch [hirsch2012differential, Section $5.2$]).
Consider an embedded Lagrangian sphere $i:L\rightarrow(M,\omega)$ where $(M,\omega)$ is a $4$-dimensional symplectic manifold. Using Weinstein’s Lagrangian neighborhood theorem, we identify a neighborhood of $i(L)$ with the cotangent bundle of $S^{2}$. Therefore the intersection number $L\cdot L$ is such that $|L\cdot L|=\chi(S^{2})\neq 0$. Now we prove that $i(L)$ is not topologically displaceable by a smooth map $g:L\rightarrow M$ homotopic to the inclusion $i$. Let us argue by contradiction. Assume that $g(i(L))\cap i(L)=\emptyset$ and $g\simeq i$. In particular, $g$ and $i$ are transverse, and since $g(i(L))\cap i(L)=\emptyset$ we have $L\cdot L=\#(g,L)=0$, where #$\#(g,L)$ is the intersection number of the map $g$ with $L$. Therefore we obtain a contradiction. Since $L$ is not topologically displaceable, in particular it is not displaceable by a Hamiltonian diffeomorphism. ∎
Corollary 3.4.
A focus-focus fiber $F^{-1}(c)$ is nondisplaceable when its multiplicity is greater than or equal to $2$.
Therefore studying (non)displaceability of focus-focus fibers is reduced to the case $k=1$, which is the generic case, see Smirnov [smirnov2013focus]. There are also explicit examples for higher multiplicity. For the case $k=2$ see for instance De Meulenaere & Hohloch [de2021family] and Section 4.3.
3.3. Displacing focus-focus fibers
In this subsection we explain how to displace focus-focus fibers in certain semitoric systems using the methods of Subsection 3.1.
Let $(M,\omega,F)$ be a semitoric system on a $4$-dimensional symplectic manifold $(M,\omega)$ and to reduce notation assume that we have a single focus-focus value $c$. Due to Subsection 3.2 we focus on the case of multiplicity $1$.
Let $\Delta_{\vec{\epsilon}}$ be a representative of the polytope invariant for the semitoric system. Without loss of generality we henceforth assume that $\vec{\epsilon}=-1$.
Let $h$ be the height invariant for the focus-focus value and $v_{1},v_{2}$ the edges that intersect at the corner introduced by the focus-focus value in the polytope invariant, as sketched in Figure 3.1.
Figure 3.1. Part of the representative of the polytope invariant for $\vec{\epsilon}=-1$. The red dot represent the focus-focus value.
Now assume that the corner introduced by the focus-focus fiber is a Delzant corner, and let $A$ be the integral affine transformation that takes it into the standard Delzant corner. By Remark 2.20 we can apply a nodal trade (which is defined in Lemma 2.15). Recall the affine distance from Equation (2.1) and denote the affine length by aff. Without loss of generality assume that $\text{aff}(v_{1})\geq\text{aff}(v_{2})$. Suppose that $h<\text{aff}(v_{2})$, and that there exists $R>0$ such that $R<\text{aff}(v_{1})$ and $0<2h<R$. Furthermore, suppose that the rectangle $[0,R]\times[0,h]$ starting at $(0,0)$ sits inside $A(\Delta_{\vec{\epsilon}})$, as is sketched in Figure 3.2. Then we obtain the following result:
Lemma 3.5.
The focus-focus fiber of the system $(M,\omega,F)$ described above is displaceable.
Proof.
Recall that for the standard Delzant corner the corner is placed at the origin. After applying a nodal trade, as defined in Lemma 2.15, the semitoric system with focus-focus fiber $F^{-1}(c)$ becomes a toric fibration $\pi:(M,\omega)\rightarrow\mathbb{R}^{2}$. Hereby the set $F^{-1}(c)$ is contained in $\pi^{-1}([0,h]\times[0,h])$. Since $h<\frac{R}{2}$, by Lemma 3.1, the set $\pi^{-1}([0,h]\times[0,h])$ is displaceable, in particular the focus-focus fiber is displaceable. ∎
Figure 3.2. Rectangle $[0,R]\times[0,h]$ sitting inside $A(\Delta_{\vec{\epsilon}})$. The red dot represents the focus-focus value.
4. (Non)displaceability in three explicit systems
In this section we consider (non)displaceability of fibers of the coupled spin oscillator, Subsection 4.1, the coupled angular momenta, Subsection 4.2, and a semitoric system on the symplectic manifold given by the octagon, Subsection 4.3.
4.1. (Non)displaceability within the coupled spin-oscillator
In this section, we study displaceability properties of fibers of the coupled spin-oscillator, a non-compact integrable system.
Let $\rho_{1},\rho_{2}>0$ be positive constants. Consider the product manifold $M=S^{2}\times\mathbb{R}^{2}$ with symplectic form $\omega=\rho_{1}\omega_{S^{2}}\bigoplus\rho_{2}\omega_{\mathbb{R}^{2}}$, where $\omega_{S^{2}}$ and $\omega_{\mathbb{R}^{2}}$ are the standard symplectic structures on the unit sphere and the Euclidean space respectively. Let $(x,y,z)$ be Cartesian coordinates on the unit sphere $S^{2}\subset\mathbb{R}^{3}$ and $(u,v)$ Cartesian coordinates on the plane $\mathbb{R}^{2}$.
Definition 4.1.
A coupled spin-oscillator is a $4$-dimensional integrable system $(M,\omega,(L,H))$, where the momentum map $F=(L,H):M\rightarrow\mathbb{R}^{2}$ is given by
(4.1) $L(x,y,z,u,v):=\rho_{1}z+\rho_{2}\frac{u^{2}+v^{2}}{2},\quad H(x,y,z,u,v):=\frac{xu+yv}{2}.$
Coupled spin-oscillators are semitoric systems, for more details and background see Pelayo & Vũ Ngǫc [pelayo2012hamiltonian]. The system has exactly one focus-focus singularity at the point $m:=(0,0,1,0,0)$. In order to understand the displaceability properties of the fibers of these systems we look at a representative of the polytope invariant, see Figure 4.1. These were computed in Pelayo & Vũ Ngǫc  [pelayo2012hamiltonian] and Alonso [alonso2019symplecticthesis, Section $5.1$].
Figure 4.1. Representative of the polytope invariant of the coupled-spin oscillator with $\vec{\epsilon}=+1$. The red dot represents the focus-focus value and is placed at $(0,0)$. Notice that the polytope is unbounded.
Theorem 4.2.
All of the fibers of a coupled spin-oscillator are displaceable.
Proof.
There are two cases:
  • First consider the fibers outside of the segment of the eigenline through the focus-focus value in Figure 4.1. By Lemma 2.22 all of these fibers are displaceable by a probe with direction $(1,0)$ starting at the facet $y=x+\rho_{1}$, possibly after getting rid of the segment of the eigenline to the focus-focus value by applying a nodal trade, see Lemma 2.15.
  • To displace the focus-focus fiber and the fibers lying on the segment of the eigenline of Figure 4.1 we use the results of Section 3.3, since one of the of the edges has infinite affine length.
Intuitively, it is the non-compactness of the $\mathbb{R}^{2}$ factor that allows us to displace every fiber.
4.2. (Non)displaceability within the coupled angular momenta
In this section, we study displaceability properties of fibers of the following systems:
Definition 4.3.
Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=S^{2}\times S^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{S^{2}}\bigoplus R_{2}\omega_{S^{2}})$ where $\omega_{S^{2}}$ is the standard symplectic form on $S^{2}$. Let $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})$ be Cartesian coordinates on $M$ induced from the ambient ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, where the smooth functions $L,H_{t}:M\rightarrow\mathbb{R}$ are given by
$\begin{cases}L(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=R_{1}z_{1}+R_{2}z_{2},\\ H_{t}(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=(1-t)z_{1}+t(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})\end{cases}$
and $R_{1}<R_{2}$. The case $R_{1}>R_{2}$ is called reverse coupled angular momenta and the case $R_{1}=R_{2}$ is called the Kepler problem.
The system has four fixed points located at $(0,0,\pm 1,0,0,\pm 1)$. Three are of elliptic-elliptic type, namely $(0,0,\pm 1,0,0,1)$, but $(0,0,1,0,0,-1)$ is of focus-focus type for $t\in\ ]t^{-},t^{+}[$ and degenerate for $t\in\{t^{-},t{+}\}$, where
$t^{\pm}:=\frac{R_{2}}{2R_{2}+R_{1}\mp 2\sqrt{R_{1}R_{2}}}.$
Note that at the parameter $t=\frac{1}{2}$ there is always a focus-focus singularity. For more details and proofs of these results, see for instance Alonso [alonso2019symplecticthesis, Section 6] and Alonso & Dullin & Hohloch [alonso2019symplectic].
4.2.1. (Non)displaceability of the fibers
We now study the (non)displaceability of the fibers of the systems $(M,\omega,F_{t})$ in Definition 4.3. Let us consider the map $\Psi:M\rightarrow M$ given by
(4.2) $\Psi(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=(-x_{1},y_{1},-z_{1},x_{2},-y_{2},-z_{2}).$
Notice that $\Psi$ is a Hamiltonian diffeomorphism since it is the time-$1$ map of a suitable normalization of the Hamiltonian $G:M\to{\mathbb{R}}$, $G(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=y_{1}+x_{2}$. Using $\Psi$, we obtain the following result:
Proposition 4.4.
Consider $(a,b)\in F_{t}(M)$ with $a\neq 0$. Then the fiber $(F_{t})^{-1}(a,b)$ is displaceable for $t\in[0,1]$.
Proof.
For $p:=(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})\in M$ we have $L(p)=R_{1}z_{1}+R_{2}z_{2}$ and
$L(\Psi(p))=-(R_{1}z_{1}+R_{2}z_{2})=-L(p).$
Therefore, whenever $a\neq 0$, we deduce
$F_{t}^{-1}(a,b)\cap\Psi(F_{t}^{-1}(a,b))=\emptyset.$
Corollary 4.5.
Let $R_{1}\neq R_{2}$ and $t\in\ ]t^{-},t^{+}[$, the focus-focus fiber of $(M,\omega,F_{t})$ is displaceable.
Proof.
The focus-focus value is $(R_{1}-R_{2},1-2t)$. Hence the result follows from Proposition 4.4. ∎
It remains to investigate the (non)displaceability properties of $F^{-1}(a,b)$ where $(a,b)\in F_{t}(M)$ with $a=0$. For the rest of this section assume $R_{2}>R_{1}$. The case $R_{1}>R_{2}$ is analogous. The case $R_{2}=R_{1}$ will be dealt with separately in Section 4.2.2.
We consider the system for $t^{-}<t<t^{+}$, i.e., there are three elliptic-elliptic singularities and one singularity of focus-focus type. For the following statement, we will make use of the polytope invariant (cf. Section 2.8) of the coupled angular momenta system. One of the representatives of the polytope invariant, after applying a horizontal translation, is drawn in Figure 4.2.
Figure 4.2. A horizontal translation of a representative of the polytope invariant for $\vec{\epsilon}=1$ of the coupled angular momenta system where $t^{-}<t<t^{+}$. The red segment stands for the cut from above to the focus-focus value along the eigenline. The orange line represents the vertical line $x=R_{2}-R_{1}$. The value $h_{1}$ is the height invariant of the system.
Proposition 4.6.
Let $R_{1}\neq R_{2}$ and $t^{-}<t<t^{+}$. Then
  • there only exists one nondisplaceable fiber in the coupled angular momenta system $F_{t}=(L,H_{t})$, namely $(F_{t}^{-1}(f^{-1}))(R_{2}-R_{1},-h_{1}+R_{1})$, where $h_{1}$ is the height invariant of the system and $f$ is the map associated with Figure 4.2, i.e, the horizontal translation of a representative of the polytope invariant (cf. Theorem 2.18). Therefore the nondisplaceable fiber is a stem and nondisplaceable by symplectomorphisms. This fiber is independent of the choice of representative of the polytope invariant.
  • All other fibers are displaceable.
  • In particular, the focus-focus fiber is displaceable.
Proof.
Note that the map $f$, defined as the composition of a horizontal translation with the straightening homeomorphism $f_{\vec{\epsilon}}$ of a representative of the polytope invariant, maps the vertical line $x=0$ in $F_{t}(M)$ to the vertical line $x=R_{2}-R_{1}$ in $f(F_{t}(M))$, see Alonso & Dullin & Hohloch [alonso2019symplectic].
We can apply Lemma 2.22 as long as the probe does not intersect the segment of the eigenline, since outside this segment the semitoric system is a toric fibration. Hence, in this polytope, by using a probe with direction $(0,\pm 1)$, every fiber of the form $F_{t}^{-1}(f^{-1}(R_{2}-R_{1},b))$ with $b\neq-h_{1}+R_{1}$ is displaceable. Furthermore, using Theorem 2.33 and Proposition 4.4, we conclude that the fiber $F_{t}^{-1}(f^{-1}(R_{2}-R_{1},-h_{1}+R_{1}))$ is nondisplaceable. In particular it is a stem, and by Theorem 2.35 nondisplaceable by symplectomorphisms. ∎
4.2.2. (Non)displaceability in the Kepler problem
Now we consider the case $R_{1}=R_{2}=:R$ which is known as the Kepler problem. Filling $R$ into the formula of $t^{\pm}$, we find $t^{-}=\frac{1}{5}$ and $t^{+}=1$.
We will study the system for $\frac{1}{5}<t<1$, where the system has precisely one focus-focus point. By Proposition 4.4, every fiber of $F_{t}$ over $(a,b)$ with $a\neq 0$ is displaceable. Let us now consider a fiber $F_{t}^{-1}(0,b)$ with $b\in\mathbb{R}$.
Proposition 4.7.
Consider the Kepler problem with $\frac{1}{5}<t<1$. Then the fiber $F_{t}^{-1}(0,b)$ is displaceable for $b\in([-1,-t[\ \cup\ \mathbb{R}^{>0})\cap F_{t}(M)$. In particular, if $\frac{1}{5}<t<\frac{1}{2}$, the focus-focus fiber $F_{t}^{-1}(0,1-2t)$ is displaceable.
Proof.
Recall the Hamiltonian diffeomorphism $\Psi$ from Equation (4.2). Consider $p:=(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})\in M$ and set $b:=H_{t}(p)=(1-t)z_{1}+t(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})$. Then
$H_{t}(\Psi(p))$ $=-(1-t)z_{1}+t(-x_{1}x_{2}-y_{1}y_{2}+z_{1}z_{2})$
$=-(1-t)z_{1}-t(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})+2tz_{1}z_{2}$
$=-H_{t}(p)+2tz_{1}z_{2}$
$=-b+2tz_{1}z_{2}.$
On $L^{-1}(0)$, we have $z_{1}=-z_{2}$. Setting $z:=z_{1}=-z_{2}$, we find that $H_{t}(\Psi(p))=-b-2tz^{2}$ on $L^{-1}(0)$. Therefore we have
$F_{t}(\Psi(F_{t}^{-1}(0,b)))\subset\{0\}\times[-b-2t,-b].$
For $b<-t$, we have $-b-2t>-t$ and hence
$\Psi(F_{t}^{-1}(0,b))\cap F_{t}^{-1}(0,b)=\emptyset.$
Moreover, for $b\in\mathbb{R}^{>0}$, we have $[-b-2t,-b]\cap\{b\}=\emptyset$, hence the desired fibers are displaceable by $\Psi$. Note that the focus-focus value is $(0,1-2t)\in{\mathbb{R}}^{2}$. Hence, for $t<\frac{1}{2}$, the focus-focus fiber $F_{t}^{-1}(0,1-2t)$ is displaceable. ∎
Let us now inquire about the (non)displaceability of the focus-focus fiber for $t\geq\frac{1}{2}$. First we need to establish some facts about the height invariant. From now on $h:=h_{t}$ is the height invariant associated with the Kepler problem $(M,\omega,F_{t})$. In Alonso & Dullin & Hohloch [alonso2019symplectic] the authors compute the height invariant of the Kepler problem and show that it is given by
$h=R\tilde{h}$
where
$\tilde{h}:=2-\frac{2}{\pi}\left(2\arctan(e^{-v})-sech(v))\right),\quad v:=\text{arctanh}\left(\frac{1-3t}{2t}\right).$
In order to understand the displaceability properties we need to understand how $h=h_{t}$ varies with $t$. Figure 4.3 shows a plot of $\tilde{h}$.
Figure 4.3. Plot of $\tilde{h}$ for $t\in\ ]t^{-},t^{+}[$ in blue. Plot of the constant line equal $1$ in yellow. Plot of the constant line equal 00 in green.
Let $t^{-}<t_{0}<t^{+}$ be the value determined by $\tilde{h}(t_{0})=1$. Define the maps $T_{0,h-2R}:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ and $A:{\mathbb{R}}^{2}\to{\mathbb{R}}^{2}$ via
(4.3) $T_{0,h-2R}(x,y):=(x,y+(h-2R))\quad\mbox{and}\quad A:=\begin{bmatrix}0&-1\\ 1&-1\end{bmatrix}\in GL(n,\mathbb{Z}).$
Proposition 4.8.
For $\frac{1}{5}<t<t_{0}$, the Kepler problem $(M,\omega,F_{t})$ has a unique nondisplaceable fiber. In particular it is a stem and nondisplaceable by symplectomorphisms. It is given by $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}((A\circ T_{0,h-2R})^{-1}(R,R)))$ where $A$ and $T_{0,h-2R}$ are the maps defined in (4.3) and $f_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the representative of the polytope invariant given in Figure 4.4. In particular, the focus-focus fiber is displaceable.
Proof.
Consider the representative with $\vec{\epsilon}=1$ of the polytope invariant plotted in Figure 4.4. Furthermore, notice that applying the transformation $A\circ T_{0,h-2R}$ to the polytope invariant in Figure 4.4 we obtain Figure 4.5. Furthermore the set $J^{-1}(0)$ is mapped to the diagonal in Figure 4.5.
Figure 4.4. Representative of the polytope invariant with $\vec{\epsilon}=1$ for the Kepler problem for $\frac{1}{5}<t<1$. The cut along the eigenline is sketched in red.
Step 1: Displaceability on the segment of the eigenline on the focus-focus value. Due to the choice of $t_{0}$ we have that for $t^{-}<t<t_{0}$, $\tilde{h}(t)>1$ and thus $2R-h<R$. Hence we can apply Lemma 3.5 in Figure 4.5 to conclude that the focus-focus fiber and every fiber of the form $F_{t}^{-1}(0,b)$, where $(0,b)\in F_{t}(M)$ with $b>1-2t$, is displaceable.
Step 2: Displaceability outside the segment of the eigenline for values in $J^{-1}(0)$. Notice that outside the segment of the eigenline the semitoric system is a toric fibration. Therefore, use Lemma 2.22 to displace every fiber in the diagonal of $A\circ T_{0,h-2R}(f_{\vec{\epsilon}}(F_{t}(M))$ that does not lie over the red segment in Figure 4.5 and is not the fiber over $(R,R)$.
Figure 4.5. Image of the nodal trade of the polytope invariant by $A\circ T_{0,h-2R}$. The red line represents the image of the cut along the eigenline associated with the focus-focus fiber.
We cannot displace the fiber over $(R,R)$ since for every probe passing trough it, the fiber is halfway the distance of the probe.
The above results combined with Proposition 4.4 show that all fibers except the fiber over $(R,R)$ in Figure 4.5 are displaceable. Theorem 2.33 requires the existence of at least one nondisplaceable fiber. Therefore the fiber over $(R,R)$ in Figure 4.5 is nondisplaceable. In particular it is a stem, and nondisplaceable by symplectomorphisms. ∎
We are left with understanding the cases $t\geq t_{0}$. In Theorem 4.12 we will obtain a complete description of the (non)displaceability properties of the Kepler problem for $t=t_{0}$: there exists only one nondisplaceable fiber, and it is given by the focus-focus fiber, in particular it is a stem and hence nondisplaceable by symplectomorphisms. For $t>t_{0}$ we will see that there exist an infinite amount of nondisplaceable fibers, noticing that these fibers are equivalent to fibers already studied in the literature, see Auroux [auroux2007mirror]. In particular the focus-focus fiber is nondisplaceable. This is done in Section 4.2.3.
We start with the following displaceability result:
Proposition 4.9.
Consider the Kepler problem for $t_{0}\leq t<1$. Let $(c_{0},c_{0})$ denote the image of focus-focus value by $A\circ T_{0,h-2R}\circ f_{\vec{\epsilon}}$, where $A$ and $T_{0,h-2R}$ are the maps defined in (4.3) and $f_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the representative of the polytope invariant given in Figure 4.4. The fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}((A\circ T_{0,h-2R})^{-1}(c,c)))$, with $c>c_{0}$ are displaceable, i.e., the fibers below the focus-focus value on the line $x=0$ in $F_{t}(M)$ are displaceable.
Proof.
The proof is analogous to Step $2$ in the proof of Proposition 4.8, it is an application of the method of probes. ∎
In order to obtain further displaceability results we need to look at another representative of the polytope invariant for the Kepler problem. We use the one given by Figure 4.6, for more details see Alonso & Dullin & Hohloch [alonso2019symplectic].
Figure 4.6. Representative of the polytope invariant with $\vec{\epsilon}=-1$ for the Kepler problem for $\frac{1}{5}<t<1$. The focus-focus values is represented in red. The cut along the eigenline is sketched in red.
Proposition 4.10.
Consider the Kepler problem for $t_{0}\leq t<1$. Let $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then the fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ with $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$ of the Kepler problem are displaceable.
Proof.
The idea of the proof is to use Lemma 2.22 on the polytope invariant given by Figure 4.6, making sure that the probe does not intersect the segment of the eigenline associated with the focus-focus fiber. We can use Lemma 2.22 since outside the segment of the eigenline the semitoric system is toric. For this purpose we note that due the choice of $t$ we have $h\leq R$.
Let $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$. Notice that the vector $v:=(1,0)$ is integrally transverse to the facet given by $y=x+(2R-h)$. Therefore we can construct the probe $(x+k,y)$ where $y=x+(2R-h)>0$. The affine distance of this probe is $2R-\frac{(y+h)}{2}$ since it will end at the facet given by $y=2x-h$. Due to Lemma 2.22 the $(0,y)$ such that $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>\frac{4}{3}R-h$ give displaceable fibers, since they lie less than halfway along the probe. ∎
We can actually improve the results of Proposition 4.10 by using nodal slides:
Proposition 4.11.
Consider the Kepler problem for $t_{0}\leq t<1$. Let $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then the fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ with $(0,y)\in f_{\vec{\epsilon}}(F_{t}(M))$ and $y>R-h$ of the Kepler problem are displaceable.
Proof.
The fibers we consider here correspond to fibers that sit above the segment of the eigenline $(x,x)$ in Figure 4.5 with $x<2R-h$. In the literature these are known as Chekanov-type tori, see Auroux [auroux2007mirror].
By applying a nodal slide, see Proposition 2.14, we can make the segment of the eigenline of Figure 4.6 as small as possible while preserving the fiber we are considering. This corresponds to considering the fiber $(x,x)$ in the polytope invariant, given by Figure 4.6, of the system for a $\tilde{t}>t$. Therefore the fiber over $(x,x)$ is displaceable by the probe that starts in the facet $y=x+(2R-h)$ with direction $(0,-1)$. Hence we obtain our desired result. ∎
We can now combine all of the previous results together to obtain a complete description of the (non)displaceability of the fibers of the Kepler problem $(M,\omega,F_{t_{0}})$.
Theorem 4.12.
Consider the Kepler problem $(M,\omega,F_{t_{0}})$. The unique nondisplaceable fiber is the focus-focus fiber. All other fibers are displaceable. The focus-focus fiber is a stem and nondisplaceable by symplectomorphisms.
Proof.
The idea of the proof is to combine the previous results to obtain that every fiber other than the focus-focus fiber is displaceable. Indeed now in detail:
  • By Proposition 4.4 every fiber of the form $F_{t_{0}}^{-1}(a,b)$ with $a\neq 0$ is displaceable.
  • By Proposition 4.9 every fiber of $F$ below the focus-focus value on the line $x=0$ is displaceable.
  • Recall that due to the definition of $t_{0}$, we have $R=h$. By Proposition 4.11 every fiber of $F$ above the focus-focus value on the line $x=0$ is displaceable.
Therefore the focus-focus fiber is a stem, and hence nondisplaceable by symplectomorphisms, see Theorem 2.35. ∎
Remark 4.13.
Since the fiber for the Kepler problem $(M,\omega,F_{t_{0}})$ is actually monotone, one could use the methods of Hong & Kim & Lau [hong2023immersed] or Rizzel & Ekholm & Tonkonog [rizell2018refined] to obtain the potential function for the focus-focus fiber, and then obtain a nondisplaceability result by computing the critical points of the potential function. However the result proven in Theorem 4.12 is stronger since we prove that the focus-focus fiber is a stem.
4.2.3. The case $t>t_{0}$ for the Kepler problem.
In this subsection we show that if $t>t_{0}$ the number of nondisplaceable fibers for the Kepler problem is infinite. In particular the focus-focus fiber of the systems is nondisplaceable. First we start with a weaker result:
Proposition 4.14.
Let $N\in\mathbb{N}$ and $R>0$. Then, for $11-\frac{1}{2^{N+2}}<t<$ 1, the Kepler problem $(M,\omega,F_{t}=(L,H_{t}))$ has at least $12^{N-1}+$ 1 nondisplaceable fibers.
To prove this result we will make use of the notions of symplectic quasi-states, superheavy, heavy, and pseudoheavy fibers summarized in Section 2.13. First note:
Lemma 4.15.
Let $(M,\omega)$ be a closed symplectic manifold and $\zeta$ a partial symplectic quasi-state in $(M,\omega)$. Let $R\neq 0$. Then $\zeta$ is a partial symplectic quasi-state for $(M,R\omega)$.
Proof.
This follows from the definition of a partial symplectic quasi-state and the fact that, if $X$ is a vector field satisfying $R\omega(X,\cdot)=dH$ for some $H\in C^{\infty}(M,{\mathbb{R}})$, then $\omega(X,\cdot)=d\left(\frac{H}{R}\right)$, i.e., $X$ is the Hamiltonian vector field of $\frac{H}{R}$ w.r.t. $\omega$. ∎
Now consider the symplectic manifold $(S^{2}\times S^{2},\omega_{S^{2}}\oplus\omega_{S^{2}})$ with Cartesian coordinates induced from ${\mathbb{R}}^{3}\times{\mathbb{R}}^{3}$, and the Hamiltonians $G:=(J,H):S^{2}\times S^{2}\rightarrow\mathbb{R}^{2}$, where
$\begin{cases}L(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=z_{1}+z_{2},\\ H(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2}):=H_{1}=x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2}.\end{cases}$
Then $(S^{2}\times S^{2},\omega_{S^{2}}\oplus\omega_{S^{2}},G)$ is the Kepler problem for $R=1,t=1$ with a change of sign in the symplectic form. Set $S_{c}:=L^{-1}(0)\cap H^{-1}(c)=G^{-1}(0,c)$. Entov & Polterovich [entov2009rigid] constructed a partial symplectic quasi-state $\zeta_{-1}$ such that $S_{-1}$ is $\zeta_{-1}$-superheavy. Furthermore, according to Fukaya & Oh & Ohta & Ono [fukaya2011spectral], for every $c\in\ ]-1,-\frac{1}{2}]$, there exists a partial symplectic quasi-state $\zeta_{c}$ such that $S_{c}$ is $\zeta_{c}$-superheavy.
By Lemma 4.15, if $R\neq 0$, then all $\zeta_{c}$ with $c\in\left[-1,-\frac{1}{2}\right]$ are symplectic quasi-states for $(S^{2}\times S^{2},R(\omega_{S^{2}}\oplus\omega_{S^{2}}))$.
Proof of Proposition 4.14.
Consider such an $S_{c}$ which is superheavy with respect to $\zeta_{c}$ as introduced above. Let $N\in{\mathbb{N}}$ and set
$A:=\left\{\left.-\frac{1}{2}-\frac{k}{2^{N}}\right|k\in\{0,\dots,2^{N-1}\}\right\}.$
By Theorem 2.31, there exists, for all $c\in A$, a pseudoheavy fiber $F_{t}^{-1}(p_{c})$ for some $p_{c}\in F_{t}(S^{2}\times S^{2})$ for the integrable systems $(S^{2}\times S^{2},R(\omega_{S^{2}}\otimes\omega_{S^{2}}),F_{t})$ with $0\leq t\leq 1$ and $R\neq 0$. In particular, by Theorem 2.30, a pseudoheavy fiber has to intersect the corresponding superheavy fiber, i.e., $F^{-1}_{t}(p_{c})\cap S_{c}\neq\emptyset$. Furthermore the set $F^{-1}_{t}(p_{c})$ is nondisplaceable. Now we have to ensure that the sets $F^{-1}_{t}(p_{c})$ are disjoint for different values of $c$. First notice that
$\qquad|H_{t}(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})-H(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})|$
$\quad=|(1-t)z_{1}+t(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})-(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})|$
$\quad=|(1-t)z_{1}-(1-t)(x_{1}x_{2}+y_{1}y_{2}+z_{1}z_{2})|$
$\quad\leq(1-t)4.$
Therefore,
$H_{t}(S_{c})\subset[c-(1-t)4,\ c+(1-t)4].$
Hence, if $t>1-\frac{1}{2^{N+2}}$, for $c\in A$ the sets $[c-(1-t)4,\ c+(1-t)4]$ do not intersect each other. Therefore the nondisplaceable fibers $F_{t}^{-1}(p_{c})$ are disjoint. Hence we obtain the desired result. ∎
Proposition 4.14 tell us that the number of nondisplaceable fibers is increasing as $t\rightarrow 1$. In Lemma 4.16 we will improve this result by noticing that when $t>t_{0}$ the number of nondisplaceable fibers is actually uncountable and in particular the focus-focus fiber of the system is nondisplaceable.
Lemma 4.16.
Let $1>t>t_{0}$ and $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then the Lagrangian torus fibers $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ of $(M,\omega,F_{t})$ with $R-h\geq y>0$ are nondisplaceable. Furthermore, the focus-focus fiber of $(M,\omega,F_{t})$ is nondisplaceable.
Proof.
The Lagrangian torus fibers under consideration are Chekanov-type tori, see [auroux2007mirror]. Furthermore due to the condition on $y$ the nondisplaceability follows from Vianna & Tonkonog [tonkonog2015low, Section $3$] or Auroux [auroux2007mirror], where the potential function of these tori is computed. For this choice of $y$ the potential function has a critical point, introducing a so called bulk deformation if necessary. For more details see for example Fukaya & Oh & Ohta & Ono [fukaya2012toric]. Therefore these tori are nondisplaceable. The focus-focus fiber is nondisplaceble since the above mentioned tori are nondisplaceable and since displaceability is an open property. ∎
Remark 4.17.
Notice that we can use the results of Lemma 4.16 to show that for $t=t_{0}$ the focus-focus fiber of the system $(M,\omega,F_{t})$ is nondisplaceable. However the result we otained in Theorem 4.12 is stronger since there we actually proved that the focus-focus fiber is a stem and hence nondisplaceable by symplectomorphisms.
We can combine all the previous results to obtain a complete description of the (non)displaceability properties of the Kepler problem $(M,\omega,F_{t})$ for $t>t_{0}$:
Theorem 4.18.
Let $t_{0}<t<1$ and $f_{\vec{\epsilon}}$ be the straightening homeomorphism associated with the polytope invariant given by Figure 4.6. Then
  • For $(x,y)\in F_{t}(M)$ with $x\neq 0$ the fiber $F_{t}^{-1}(x,y)$ is displaceable.
  • For $R-h<y\leq 2R-h$ the fiber $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ is displaceable.
  • For $0<y\leq R-h$ the fiber $F_{t}^{-1}(f_{\vec{\epsilon}}^{-1}(0,y))$ is nondisplaceable.
  • The focus-focus fiber is nondisplaceable.
  • For $(x,y)\in F_{t}(M)$ and $y<1-2t$ the fiber $F_{t}^{-1}(0,y)$ is displaceable.
The content of Theorem 4.18 is summarized in Picture 4.7.
Figure 4.7. Representative of the polytope invariant with $\vec{\epsilon}=-1$ for the Kepler problem for $t_{0}<t<1$. The cut along the eigenline is sketched in red. The nondisplaceable fibers are identified in green, all the other fibers are displaceable.
In the following lemma we obtain a simple result about the nonpseudoheaviness of the focus-focus fiber for the symplectic quasi-states $\zeta_{a}$ as $t\rightarrow 1$.
Lemma 4.19.
Let $a\in\ ]-1,-\frac{1}{2}]$. For $\frac{2}{3+a}<t<1$ the focus-focus fiber of the semitoric system $(M,\omega,F_{t})$ is not pseudoheavy with respect to $\zeta_{a}$.
Proof.
By Theorem 2.30 if the focus-focus fiber is $\zeta_{a}$-pseudoheavy then it intersects the superheavy fiber $S_{a}$. Therefore, there must exist a point $(x_{1},y_{1},z_{1},x_{2},y_{2},z_{2})\in S^{2}\times S^{2}$ such that
(4.4) $1-2t=(1-t)z_{1}+ta,$
i.e., $z_{1}=\frac{1+(a-2)t}{1-t}$. If $t>\frac{2}{3+a}$ then $z_{1}<-1$, hence we obtain a contradiction and the desired result. ∎
Now we summarize the results about the (non)displaceability properties of the focus-focus fiber of the systems $(M,\omega,F_{t})$ in the following proposition:
Proposition 4.20.
Consider the Kepler problem $(M,\omega,F_{t})$ for $\frac{1}{5}<t<1$. Then for
  • $t<t_{0}$, the focus-focus fiber is displaceable.
  • $t=t_{0}$ the focus-focus fiber is a stem, hence nondisplaceable by symplectomorphisms.
  • $t>t_{0}$, the focus-focus fiber is nondisplaceable.
4.3. A semitoric system induced by the octagon
Consider the toric symplectic manifold $(M,\omega)$ induced by the octagon $\Delta$ with vertices
$\{(1,0),(0,1),(2,0),(0,2),(1,3),(2,3),(3,1),(3,2)\}.$
In De Meulenaere & Hohloch [de2021family, Section $3.1$] it is shown that $M=\tilde{L}^{-1}(0)/N$, where
(4.5) $\tilde{L}^{-1}(0):=\left\{(z_{1},...,z_{8})\in\mathbb{C}^{8}\left|\ \begin{aligned} &|z_{1}|^{2}+|z_{5}|^{2}=6\\ &|z_{2}|^{2}+|z_{5}|^{2}+|z_{7}|^{2}=10\\ &|z_{3}|^{2}+|z_{7}|^{2}=6\\ &|z_{4}|^{2}-|z_{5}|^{2}+|z_{7}|^{2}=4\\ &|z_{5}|^{2}-|z_{6}|^{2}+|z_{7}|^{2}=2\\ &|z_{5}|^{2}-|z_{7}|^{2}+|z_{8}|^{2}=4\end{aligned}\right.\right\}$
and $N\simeq\mathbb{T}^{6}$ acts on $\mathbb{C}^{8}$ as
(4.6) $(t,z)\mapsto$
$(e^{it_{1}}z_{1},e^{it_{2}}z_{2},e^{it_{3}}z_{3},e^{it_{4}}z_{4},e^{i(t_{1}+t_{2}-t_{4}-t_{5}+t_{6})}z_{5},e^{it_{5}}z_{6},e^{i(t_{2}+t_{3}+t_{4}-t_{5}-t_{6})}z_{7},e^{it_{6}}z_{8}).$
Points on $(M,\omega)$ are written as equivalence classes of the form $[z]=[z_{1},...,z_{8}]$ with $z_{k}=x_{k}+iy_{k}\in\mathbb{C}$ for $1\leq k\leq 8$. The momentum map of the toric system $(M,\omega)$ is given as follows:
Theorem 4.21.
(De Meulenaere &\&& Hohloch, [de2021family, Theorem 1.1]) Let $F=(J,H):M\rightarrow\mathbb{R}^{2}$ be given by
$\begin{cases}J([z_{1},...,z_{8}])=\frac{1}{2}|z_{1}|^{2},\\ H([z_{1},...,z_{8}])=\frac{1}{2}|z_{3}|^{2}.\end{cases}$
Then $F$ is a momentum map of an effective Hamiltonian $2$-torus action satisfying $F(M)=\Delta$. In particular, $F$ has eight elliptic-elliptic singular points.
An integrable system $(M,\omega,F)$ is of toric type if $F:M\rightarrow\mathbb{R}^{2}$ is proper and if there exists an effective, Hamiltonian $\mathbb{T}^{2}$-action on $M$ whose momentum map is of the form $f\circ F$ where $f:F(M)\rightarrow f(F(M))$ is a diffeomorphism. After perturbing the integral $H$ by $X:=\mathfrak{R}(Z)$, where $Z[z_{1},...,z_{8}]:=\overline{z_{2}z_{3}z_{4}}z_{6}z_{7}z_{8}$ and $\mathfrak{R}$ stands for the real part of the function, one is able to obtain a family of semitoric systems on $(M,\omega)$.
Theorem 4.22.
(De Meulenaere &\&& Hohloch, [de2021family, Theorem 1.2]) Set $F_{t}:=(J,H_{t}):=(J,(1-2t)H+t\gamma X):(M,\omega)\rightarrow\mathbb{R}^{2}$. Then, $(M,\omega,F_{t})_{0\leq t\leq 1}$ is toric for $t=0$, of toric type for $0<t<t^{-},$ semitoric for $t^{-}<t<t^{+}$, and again of toric type for $t^{+}<t\leq 1$ where
$10<t^{-}:=\frac{1}{2(1+24\gamma)}<\frac{1}{2}<t^{+}:=\frac{1}{2(1-24\gamma)}<$ 1 .
For all $t\in[0,1]$, the system $F_{t}$ has precisely eight fixed-points of which four are always elliptic-elliptic. The other four pass at $t=t^{-}$ from elliptic-elliptic via a Hamiltonian-Hopf bifurcation to focus-focus. At $t=t^{+}$, these four focus-focus points turn again back into elliptic-elliptic via a Hamiltonian-Hopf bifurcation.
In more detail:
Proposition 4.23.
(De Meulenaere &\&& Hohloch, [de2021family, Proposition 3.7]) The focus-focus points of the semitoric systems $(M,\omega,F_{t})$, for $t^{-}<t<t^{+}$ are
  • $B:=[\sqrt{2},0,0,\sqrt{2},2,2\sqrt{2},\sqrt{6},\sqrt{6}]$;
  • $D:=[2,\sqrt{2},0,0,\sqrt{2},\sqrt{6},\sqrt{6},2\sqrt{2}]$;
  • $C:=[2,2\sqrt{2},\sqrt{6},\sqrt{6},\sqrt{2},0,0,\sqrt{2}]$;
  • $A:=[\sqrt{2},\sqrt{6},\sqrt{6},2\sqrt{2},2,\sqrt{2},0,0].$
Furthermore $A,B\in J^{-1}(1)$ and $C,D\in J^{-1}(2)$.
Proposition 4.24.
(De Meulenaere &\&& Hohloch, [de2021family, Proposition 1.3]) At $t=\frac{1}{2}$, the system $F_{\frac{1}{2}}$ has precisely two focus-focus fibers, each of which contains precisely two focus-focus points so that each of these two fibers has the shape of a double pinched torus.
Our goal is to study the symplectic topology of the fibers of these semitoric systems, in particular of the focus-focus fibers. By Proposition 3.3, for $t=\frac{1}{2}$ the focus-focus fibers of the semitoric system are nondisplaceable, due to the presence of Lagrangian spheres.
Lemma 4.25.
There exist symplectomorphisms of $(M,\omega)$ that interchange the focus-focus fibers of the semitoric system $(M,\omega,F_{t})$ for each $t^{-}<t<t^{+}$.
Proof.
Consider the symplectomorphisms
$\Psi_{1}(z_{1},...,z_{8})=(z_{1},z_{8},z_{7},z_{6},z_{5},z_{4},z_{3},-z_{2}),$
$\Psi_{2}(z_{1},...,z_{8})=(z_{5},z_{4},z_{3},z_{2},z_{1},z_{8},z_{7},z_{6})$
of $(\mathbb{C}^{8},\omega_{0})$. Recall that $(M,\omega)$ is obtained as symplectic reduction of $(\mathbb{C}^{8},\omega_{0})$ on a level set $\tilde{L}^{-1}(0)$, see Equation (4.5), by a torus action of $\mathbb{T}^{6}$, see Equation (4.6). Note that $\Psi_{1}$ and $\Psi_{2}$ preserve $\tilde{L}^{-1}(0)$ and are invariant under the action, hence they define symplectomorphisms on $(M,\omega)$. The result follows from considering the projections of $\Psi_{1},\Psi_{2}$,$\Psi_{1}\circ\Psi_{2}$ and $\Psi_{2}\circ\Psi_{1}$ to $(M,\omega)$. The symplectomorphisms do not depend on $t$. ∎
Since (non)displaceability is invariant under symplectomorphism it is enough to consider one of the focus-focus fibers.
4.3.1. Symplectic topology of the fibers of the toric system
Before studying the symplectic topology of the fibers of the semitoric systems, we first need to understand the symplectic topology of the torus fibers given by the octagon $\Delta$.
Lemma 4.26.
Every fiber except those over the points $\{(1,1),(1,2),(\frac{3}{2},\frac{3}{2}),(2,1),(2,2)\}$ in the toric system corresponding to $\Delta$ is displaceable by probes.
Proof.
Application of Lemma 2.22: depending on the value one can use the probe with direction (±$(\pm 1,0)$ or $(0,\pm 1)$, see Figure 4.8. ∎
Figure 4.8. The fibers in the octagon that aren’t displaceable by probes are highlighted at red.
Now we study the (non)displaceability of the fibers of $\{(1,1),(1,2),(\frac{3}{2},\frac{3}{2}),(2,1),(2,2)\}$.
Theorem 4.27.
Let $(M,\omega)$ be the symplectic toric manifold given by $\Delta$. Let $\{p_{1},...,p_{5}\}=\{(1,1),(2,1),(\frac{3}{2},\frac{3}{2}),(1,2),(2,2)\}$. Then there exist $5$ different symplectic quasi-states $\zeta_{1},...,\zeta_{5}$ for $(M,\omega)$ such that for each $p_{i},\ i=1,..,5,$ the fiber $L_{p_{i}}$ over the point $p_{i}$ is superheavy with respect to $\zeta_{i}$. Therefore, the fibers $L_{i}$ are nondisplaceble.
Proof.
We present the argument for the fiber over $(\frac{3}{2},\frac{3}{2})$, the other ones are analogous. See Borman [borman2013quasi] and Abreu & Macarini [abreu2013remarks] for more details on the general method.
  • Notice that $(M,\omega)$ can be viewed as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{C}\mathbb{P}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$ for an appropriate choice of subtorus with associated Lie algebra vectors and appropriate chosen symplectic volume of $\mathbb{CP}^{2}$ and $\mathbb{CP}^{1}$, see Figure 4.9. For more details see Section 2.10 and Abreu & Macarini [abreu2013remarks].
  • Notice that the central fiber in $(\mathbb{CP}^{1},\omega_{0})$ and the fiber over $(\frac{1}{3},\frac{1}{3})$ in $(\mathbb{CP}^{2},\omega_{FS})$ are stems, and in particular superheavy for any quasi-state, see Polterovich & Rosen [polterovich2014function, Proposition $6.1.13$].
  • For each $\mathbb{CP}^{1}$ and for $\mathbb{CP}^{2}$ we choose quasi-states as in Borman [borman2013quasi, Theorem $3.1$] to create a product quasi-state in $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$, see Borman [borman2013quasi, Corollary $3.2$], such that the product of the above mentioned fibers is superheavy with respect to this quasi state. Furthermore, this quasi-state satisfies the PB-inequality.
  • Now apply Theorem 2.36 to obtain the desired result. The symplectic volume of $\mathbb{CP}^{1}$ and $\mathbb{CP}^{2}$ are chosen such that the superheavy fiber obtained as the product of superheavy fibers projects onto the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in $\Delta$.
Figure 4.9. $(M,\omega)$ as the symplectic reduction of $\mathbb{CP}^{2}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}\times\mathbb{CP}^{1}$. The dotted lines come in pairs, and each correspond to a reduction by $\mathbb{CP}^{1}$. The triangle corresponds to $\mathbb{CP}^{2}$.
4.3.2. Symplectic topology of the fibers of the semitoric systems
Now that we understand the rigidity properties of the fibers of the toric system given by the octagon we are ready to study the semitoric case. When there are $4$ focus-focus values, for an appropriate choice of $\vec{\epsilon}$, the polytope invariant of the semitoric systems is given by the octagon, see Figure 4.10.
Figure 4.10. Representative of the polytope invariant for the semitoric system when $\vec{\epsilon}=(-1,1,-1,1)$. The dotted lines represent the eigenrays associated with the focus-focus fibers. The red circles represent the $4$ different focus-focus values.
Throughout the text we use the polytope invariant of the semitoric for different choices of $\vec{\epsilon}$. We refer to Appendix A for a detailed computation of these representatives of the polytope invariant.
Lemma 4.28.
Consider the semitoric system given by $(M,\omega,F_{t}=(J,H_{t}))$ for $t^{-}<t<t^{+}$. Then there exist nondisplaceable fibers $F_{1},F_{2},F_{3}$ such that $F_{1}\in J^{-1}(1)$, $F_{2}\in J^{-1}(2)$ and $F_{3}\in J^{-1}(\frac{3}{2})$.
Proof.
We prove that $F_{1}$ exists and that $F_{1}\in J^{-1}(1)$. The proof for $F_{2}$ and $F_{3}$ is analogous.
Recall that for the toric system given by the octagon we have two superheavy fibers that lie inside $J^{-1}(1)$, namely $L_{(1,1)}$ and $L_{(1,2)}$ with respect to the symplectic quasi-states $\zeta_{(1,1)}$ and $\zeta_{(1,2)}$, see Theorem 4.27.
Using Theorem 2.31 we obtain that for the semitoric systems given by $F_{t}$ there exist pseudoheasvy fibers $F^{-1}_{t}(y_{(1,1)})$ and $F_{t}^{-1}(y_{(1,2)})$, for some $y_{(1,1)},y_{(1,2)}\in\mathbb{R}^{2}$, with respect to $\zeta_{(1,1)}$ and $\zeta_{(1,2)}$. By Theorem 2.30 both fibers lie inside $J^{-1}(1)$, since they must intersect the corresponding superheavy fibers of the corresponding quasi-state. Therefore a nondisplaceable fiber for the semitoric system $(M,\omega,F_{t})$ must exist inside $J^{-1}(1)$. ∎
Henceforth, throughout this subsection, we focus on the semitoric systems $(M,\omega,F_{t})$ that have $4$ focus-focus values.
Lemma 4.29.
Consider the semitoric system $(M,\omega,F_{t})$, where $t^{-}<t<t^{+}$. All the fibers corresponding to points on the polytope invariant given by Figure 4.10 that do not lie on the eigenrays and are not the points $\{(1,1),(2,1),(1,2),(2,2),(\frac{3}{2},\frac{3}{2})\}$ are displaceable.
Proof.
Application of the method of probes (Lemma 2.22). ∎
Lemma 4.30.
Consider the semitoric system $(M,\omega,F_{t})$, where $t^{-}<t<t^{+}$. The torus fiber corresponding to the point $(\frac{3}{2},\frac{3}{2})$ on a representative of the polytope invariant of the system $(M,\omega,F_{t})$, where $t^{-}<t<t^{+}$, is nondisplaceable.
Proof.
First note that the fiber does not depend on the choice of representative of the polytope invariant since the eingenlines associated with the focus-focus values do not intersect the vertical line $x=\frac{3}{2}$. See Figure 4.10 for a representative of the polytope invariant in the case $\vec{\epsilon}=(-1,1,-1,1)$. Applying Lemma 2.22 in a representative of the polytope invariant, we can displace every fiber in $J^{-1}(\frac{3}{2})$ that is not the fiber over the point $(\frac{3}{2},\frac{3}{2})$. Then use Lemma 4.28.
Alternatively, we could do a nodal trade ( see Lemma 2.15) on the representative of the polytope invariant for $\vec{\epsilon}=(-1,1,-1,1)$ and use Proposition 4.27. ∎
Lemma 4.31.
Consider the semitoric system $(M,\omega,F_{t})$, where $t^{-}<t<t^{+}$. Suppose that the eigenlines associated with the focus-focus fibers in the representative of the polytope invariant given by Figure 4.10 do not intersect the points $\{(1,1),(1,2),(2,1),(2,2)\}$. Then the torus fibers corresponding to the points
$\{(1,1),(1,2),(2,1),(2,2)\}$
are nondisplaceable.
Proof.
Apply a nodal trade (see Lemma 2.15) not changing the fibration type over the considered values, and use Proposition 4.27. ∎
Let us try to understand the rigidity properties of the focus-focus fibers and of the fibers that lie over the eigenrays in Figure 4.10. We will consider the focus-focus fiber $F_{0}$ in $J^{-1}(1)$ with lower $H_{t}$ value. Due to Lemma 4.25 it is enough to understand the (non)displaceability properties of $F_{0}$ in order to understand the (non)displaceability properties of all focus-focus fibers of the system. Let $h$ be the height invariant associated with the focus-focus fiber $F_{0}$. We will focus on the cases $0<h\leq\frac{3}{2}$, the case $h>\frac{3}{2}$ is analogous.
4.3.3. Case $10<h<$ 1
In this section we show that for $h<1$ the focus-focus fiber is displaceable. To this aim consider the polytope invariant which has a representative given by Figure 4.11.
Figure 4.11. Representative of the polytope invariant for $\vec{\epsilon}=(-1,1,1,1)$. The dotted lines represent the eigenrays associated with the focus-focus values. The red dots represent the focus-focus values.
After applying a suitable integral affine transformation, we obtain the polytope in Figure 4.12.
Figure 4.12. Image after applying a suitable integral affine transformation to a representative for the polytope invariant with $\vec{\epsilon}=(-1,1,1,1)$ and $h<1$. The dotted lines represent the eigenrays associated with the focus-focus values. The red dots represent the focus-focus values.
Lemma 4.32.
If $h<1$ then the focus-focus fiber $F_{0}$ of the semitoric system is displaceable. Furthermore, every fiber in the eigenrays of the focus-focus fibers of Figure 4.11 is displaceable.
Proof.
Apply a big enough nodal slide on the eigenrays of the focus-focus fibers of Figure 4.12 that are not $F_{0}$ so that Lemma 3.5 ensures that the focus-focus fiber is displaceable. Notice that the way we displace the focus-focus fiber $F_{0}$ also displaces the fibers that lie on the eigenray associated with the focus-focus fiber $F_{0}$ in Figure 4.11. An analogous argument applies to the fibers on the eigenrays of the other focus-focus fibers.
Alternatively, to displace these torus fibers, one could also use Lemma 2.22, on the representative of the polytope invariant for $\vec{\epsilon}=(1,1,1,1)$ and $\vec{\epsilon}=(-1,-1,-1,-1)$. ∎
We can summarize the previous discussion in the following corollary:
Corollary 4.33.
[…]
4.3.4. Case $h\geq 1$
Let us consider the height invariant greater than or equal to $1$. Figure 4.13 shows a representative of the polytope invariant in the case $1<h<\frac{3}{2}$. Notice that in the case $h=\frac{3}{2}$ we have $2$ focus-focus values instead of $4$.
Figure 4.13. Polytope invariant for $\vec{\epsilon}=(-1,1,-1,1)$ with height invariant $h\geq 1$. The dotted horizontal lines indicate how big the eigenrays must be. The dotted vertical lines illustrate the eigenrays associated with the focus-focus values. The red dots represent the focus-focus values.
4.3.5. The case $h=1$
We show that the focus-focus fibers are nondisplaceable. Recall that by Lemma 4.28 a nondisplaceable fiber exists in the level set $J^{-1}(1)$.
Using polytopes for different choices of $\vec{\epsilon}$ and Lemma 2.22 we will prove that every fiber of $F_{t}$ lying inside $J^{-1}(1)$ that is not one of the focus-focus fibers must be displaceable.
Lemma 4.34.
Consider the system $(M,\omega,F_{t})$, where $t$ is such that $h=1$. Then every fiber in $J^{-1}(1)$ and in $J^{-1}(2)$ that is not a focus-focus fiber is displaceable.
Proof.
We only consider the level set $J^{-1}(1)$ since the proof for the level set $J^{-1}(2)$ is analogous. There are $3$ cases to consider in order to displace the desired fibers:
  • Applying Lemma 2.22 to Figure 4.14 we conclude that every fiber of the form $(f_{\vec{\epsilon}}\circ F_{t})^{-1}(1,b)$ with $b<1$ is displaceable, where $f_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the polytope invariant of Figure 4.14.
  • Applying Lemma 2.22 to Figure 4.15 we conclude that every fiber of the form $(g_{\vec{\epsilon}}\circ F_{t})^{-1}(1,b)$ with $b>2$ is displaceable, where $g_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the polytope of Figure 4.15.
  • Applying Lemma 2.22 to Figure 4.16 2.22 we conclude that every fiber of the form $(h_{\vec{\epsilon}}\circ F_{t})^{-1}(1,b)$ with $21<b<$ 2, where $h_{\vec{\epsilon}}$ is the straightening homeomorphism associated with the polytope of Figure 4.16, is displaceable.
Figure 4.14. Representative of the polytope invariant for $\vec{\epsilon}=(1,1,1,1)$ and $h=1$. The dotted lines are the eigenrays associated with the focus-focus values. The red dots are the focus-focus values.
Figure 4.15. Representative of the polytope invariant for $\vec{\epsilon}=(-1,-1,-1,-1)$ and $h=1$. The dotted lines are the eigenrays associated with the focus-focus values. The red dots are the focus-focus values.
Figure 4.16. Representative for $\vec{\epsilon}=(-1,1,-1,1)$ and $h=1$. The dotted lines are the eigenrays associated with the focus-focus values. The red dots are the focus-focus values.
Corollary 4.35.
Consider the semitoric system $(M,\omega,F_{t})$, where $t$ is such that $h=1$. Then the focus-focus fibers are nondisplaceable.
Proof.
Combining Lemma 4.34 and Lemma 4.28 we obtain that one of the focus-focus fibers in $J^{-1}(1)$ is nondisplaceable. Due to Lemma 4.25 all focus-focus fibers are nondisplaceable. ∎
Corollary 4.36.
Consider the semitoric system $(M,\omega,F_{t})$ where $t$ is such that $h=1$. Then the focus-focus fibers and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in a representative of the polytope invariant are nondisplaceable. Every other fiber is displaceable.
Proof.
Combine Lemma 4.29 with Lemma 4.34, Lemma 4.30, and Corollary 4.35. ∎
4.3.6. The case $\frac{3}{2}\geq h>1$
This case is analogous to the previous one.
Lemma 4.37.
Consider the system $(M,\omega,F_{t})$, where $t$ is such that $\frac{3}{2}\geq h>1$. Then every fiber in $J^{-1}(1)$ and in $J^{-1}(2)$ that is not a focus-focus fiber is displaceable.
Proof.
Let $h<\frac{3}{2}$. The proof is analogous to the proof of Lemma 4.34 with the addition that one needs to do certain nodal slides, see Proposition 2.14, on the polytope invariants of $\vec{\epsilon}=(1,1,1,1)$ and $\vec{\epsilon}=(-1,-1,-1,-1)$ in order to obtain probes with the necessary length. If $h=\frac{3}{2}$ use the polytope invariant of $\vec{\epsilon}=(1,1)$ and $\vec{\epsilon}=(-1,-1)$. ∎
Lemma 4.38.
Consider the system $(M,\omega,F_{t})$, where $t$ is such that $h=\frac{3}{2}$. Then every fiber not in $J^{-1}(1)$, not in $J^{-1}(2)$ and not the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in a representative of the polytope invariant is displaceable.
Proof.
Analogous to the proof of Lemma 4.29. Apply the method of probes to a representative of the polytope invariant. ∎
Lemma 4.39.
Consider the system $(M,\omega,F_{t})$ where $t$ is such that $h=\frac{3}{2}$. Then the focus-focus fibers are nondisplaceable.
Proof.
The focus-focus fibers have multiplicity $2$, hence they are nondisplaceable, due to the presence of Lagrangian spheres, see Proposition 3.3. ∎
Combining all the previous results we obtain the following corollary:
Corollary 4.40.
Consider the semitoric system $(M,\omega,F_{t})$ where $t$ is such that $1\leq h\leq\frac{3}{2}$. Then the focus-focus fibers and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in a representative of the polytope invariant are nondisplaceable. Every other fiber is displaceable.
We can summarize the previous discussion in the following statement:
Corollary 4.41.
Consider the semitoric system $(M,\omega,F_{t})$ for $t^{-}<t<t^{+}$.
  • If the height invariant $h$ satisfies $1\leq h<\frac{3}{2}$ then there exist $5$ nondisplaceable fibers. The focus-focus fibers and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in a representative of the polytope invariant. All other fibers are displaceable.
  • If $h=\frac{3}{2}$ then there exist $3$ nondisplaceable fibers. The focus-focus fibers and and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in a representative of the polytope invariant. All other fibers are displaceable.
For different choices of $h$ we can summarize the (non)displaceability properties of the focus-focus fibers of the system in the following corollary:
Corollary 4.42.
The focus-focus fibers have the following (non)displaceability properties with respect to different choices of the height invariant of $F_{0}$:
  • If $h<1$ then the focus-focus fibers are displaceable;
  • If $1\leq h\leq\frac{3}{2}$ then the focus-focus fibers are nondisplaceable.
Appendix A Polytope invariant for the semitoric system induced by the octagon
In order to compute the representatives of the polytope invariant of the semitoric system of section 4.3 we first need to recall some definitions and results.
Let $(M,\omega)$ be a closed symplectic manifold of dimension $2n$. The Liouville measure on $(M,\omega)$ is the measure induced by the volume form $\frac{1}{(2\pi)^{n}}\frac{\omega^{n}}{n!}$. The Duistermaat-Heckman measure $\mu_{J}$ for a Hamiltonian $J:M\rightarrow\mathbb{R}$ inducing an $S^{1}$-action is defined as $\mu_{J}([a,b])=\text{vol}(J^{-1}([a,b]))$ where vol is taken with respect to the Liouville measure in $M$. By Duistermaat & Heckman [duistermaat1982variation]
$\mu_{J}=\rho_{J}(x)\frac{|dx|}{2\pi}$
where the density $\rho_{J}(x)$, sometimes called the Duistermaat-Heckman function, is a continuous function, equal to the symplectic volume of the reduced orbifold $J^{-1}(x)/S^{1}$.
Let $J$ be the momentum map of an effective Hamiltonian $S^{1}$-action on $(M,\omega)$. For each subgroup $G\subset S^{1}$, let $M^{G}$ be the set of points in $M$ whose stabilizer is $G$. The connected components of $M^{S^{1}}$ are symplectic submanifolds, hence either points or surfaces, see Karshon [karshon1996periodic] for more details.
Lemma A.1.
(Chaperon, [chaperon1983quelques]) Let $(M,\omega,J)$ be as above. For each $p\in M^{S^{1}}$ there exist neighborhoods $U\subset M$ of $p$, $U_{0}\subset\mathbb{C}^{2}$ of $(0,0)$, and a symplectomorphism $\Psi:(U,\omega)\rightarrow(U_{0},\omega_{0})$, where $\omega_{0}=\frac{i}{2}(dz_{1}\wedge d\overline{z}_{1}+dz_{2}\wedge d\overline{z}_{2})$ making the following diagram commute
${{(U,\omega)}}$${{(U_{0},\omega_{0})}}$${\mathbb{R}}$$\scriptstyle{J}$$\scriptstyle{\Psi}$$\scriptstyle{J_{0}}$
with $J_{0}(z_{1},z_{2})=J(p)+\frac{m_{1}}{2}|z_{1}|^{2}+\frac{m_{2}}{2}|z_{2}|^{2}$.
Since the action is effective, the integers $m_{1}$ and $m_{2}$ are relatively prime. We call the integers $m_{1}$ and $m_{2}$ the isotropy weights at $p$.
Theorem A.2.
(Vũ Ngǫc, [vu2007moment, Theorem 3.8]) Let $(M,\omega,\Phi)$ be a semitoric system and $f_{\vec{\epsilon}}$ a straightening homeomorphism associated with a representative of the polytope invariant. Furthermore, let $x\in\mathbb{R}$ be a regular value of $J$. If $\alpha^{+}(x)$ (resp. $\alpha^{-}(x)$) denotes the slope of the top (resp. bottom) boundary of the polytope $f_{\vec{\epsilon}}\circ\Phi(M)$, then the derivative of the Duistermaat-Heckman function is
$\rho_{J}^{\prime}(x)=\alpha^{+}(x)-\alpha^{-}(x)$
and is locally constant on $J(M)\setminus\{\pi_{x}(f_{\vec{\epsilon}}(\Sigma_{0}(\Phi))\}\in\mathbb{R}$, where $\Sigma_{0}(\Phi)$ is the set of critical values of $\Phi$ of maximal corank and $\pi_{x}$ is the projection $(x,y)\mapsto x$. If $(x,y)\in\Sigma_{0}(\Phi)$ then
$\rho^{\prime}_{J}(x+0)-\rho^{\prime}_{J}(x-0)=-\sum_{j}k_{j}-e^{+}-e^{-},$
where the sum runs over the set of all indices $j$ such that $\pi_{x}(c_{j})=x$ and $e^{+}$ (resp. $e^{-}$) is non-zero if and only if an elliptic top vertex (resp. a bottom vertex) projects down onto $x$. If this occurs then
$e^{\pm}=-\frac{1}{a^{\pm}b^{\pm}}\geq 0,$
where $a^{\pm},b^{\pm}$ are the isotropy weights for the $S^{1}$-action at the corresponding vertices.
With Theorem A.2, we are able to compute representatives of the polytope invariant (see Section 2.8) for the semitoric system given in Section 4.3. We focus on the case of $4$ focus-focus values $(A,B,C,D)$. Furthermore, recall that there are $4$ elliptic-elliptic values. We label them $e_{1}$, $e_{2}$, $e_{3}$, $e_{4}$. We now compute a representative of the polytope invariant for the case $\vec{\epsilon}=(1,-1,1,-1)$, the other cases are analogous.
Proposition A.3.
Let $(M,\omega,F_{t})$ for $t^{-}<t<t^{+}$ be the semitoric system defined in Section 4.3. Furthermore assume that $t\neq\frac{1}{2}$ so there are $4$ focus-focus values of $(M,\omega,F_{t})$. A representative of the polytope invariant for $\vec{\epsilon}=(-1,1,-1,1)$ of the semitoric system $(M,\omega,F_{t})$ is the octagon given by Figure 4.10.
Proof.
Consider the image of $F(M)$ under the straightening homeomorphism $f_{\vec{\epsilon}}$. By abuse of notation we identify each $e_{i}$, $i=1,..,4$, with its image under $f_{\vec{\epsilon}}$. Let $c_{i}$ for $i=1,...,4$ be the new vertices caused by the cut at the focus-focus critical values $(A,B,C,D)$, respectively. In order to obtain the representative of the polytope invariant we need to compute the slopes of the edges connecting these new introduced vertices with the vertices coming from the elliptic-elliptic values. Let $l_{1}$ be the edge connecting $e_{2}$ to $c_{1}$, $l_{2}$ connect $c_{1}$ to $c_{3}$, $l_{3}$ connect $c_{3}$ to $e_{4}$, $l_{4}$ connect $e_{4}$ to $e_{3}$, $l_{5}$ connect $c_{4}$ to $e_{3}$, $l_{6}$ connect $c_{2}$ to $c_{4}$, $l_{7}$ connect $e_{1}$ to $c_{2}$ and $l_{8}$ connect $e_{1}$ to $e_{2}$. Notice that $l_{8}$ and $l_{4}$ are vertical.
First notice that due to use of an integral affine transformation that leaves a vertical line invariant we may assume that $e_{1}=(0,1)$ and that $l_{7}$ has slope $-1$. Let sl denote the slope of an edge. Using Theorem A.2, we obtain
(A.1) $(\text{sl}(l_{3})-\text{sl}(l_{5}))-(\text{sl}(l_{2})-\text{sl}(l_{6}))$ $=-2$
(A.2) $\text{sl}(l_{1})-\text{sl}(l_{7})$ $=2$
(A.3) $-(\text{sl}(l_{3})-\text{sl}(l_{5}))$ $=2.$
Since $\text{sl}(l_{7})=-1$, we get $\text{sl}(l_{1})=1$. In order to determine the slope of the other edges we use the monodromy of the system, more specifically, using item four of Theorem 2.18 we obtain $\text{sl}(l_{2})=0$. Then using item four of Theorem 2.18 iteratively and equations (A.1), (A.2), (A.3) we obtain the desired result. ∎
[…]
Sonja Hohloch &\&& Pedro Santos
University of Antwerp
Department of Mathematics
Middelheimlaan 1
B-2020 Antwerpen, Belgium
E-mail: sonja DOT hohloch AT uantwerpen DOT be
E-mail: pedro DOT santos AT uantwerpen DOT be